A Complete Bibliography of Publications in Statistical Methodology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.10

Title word cross-reference

(X, Y) [444]. (X, Y), Z [485]. 2 [357, 11]. 2 × 2 [573, 319]. 3 [160]. ^{x} [131]. E_{fNOD} [212]. g [200]. INAR(1) [346]. L [530]. L_{2} [554]. M [613, 149, 560]. R [112]. L [530]. S [377]. n [494]. P [227, 386, 408]. P (Y < X) [385]. ϕ [573]. T [347, 128]. X [240]. X/Z [444, 485]. Y/Z [444, 485]. Z [444, 167].

-Copulas [11]. -divergence [573]. -estimation [613]. -estimators [530, 560]. -moments [530]. -optimal [212, 240]. -point [357]. -smoothers [149]. -splines [227]. -square [112]. -test [128]. -tests [167]. -Treatment [200]. -values [386, 408].

/G/1 [131].

1 [131, 139].

7 [298].

8 [393].

Abbott [148]. Abbott-adjusted [148]. absolutely [272]. absorption [6]. abundance [467, 468]. accelerated [175]. acceptance [557, 390]. accidents [89]. acoustic [357]. acquisition [147]. action [238]. activation [437]. acute [597]. Adaptive [121, 90, 534, 122, 179, 563, 358]. additive [486, 552, 540]. adequacy [431]. adjust [233, 259]. adjusted [148, 276]. Adjusting [328]. adjustment [91]. admissibility [131]. advantages [186]. advice [269]. affinity [586]. Africa [264]. against [88, 536]. age [88]. ageing [536]. aging [477]. agreement [219, 321, 384]. Akaike [423]. al. [283, 281, 280, 285, 284]. algorithm [258, 162]. algorithms [364]. allocation [122, 21, 263]. almost [77]. Alpert [377]. altered [115]. alternative [457, 421]. always [444, 295, 376]. among [602]. analogues [403]. analyses [200, 441]. Analysing [266]. Analysis [389, 571, 156, 487, 301, 108, 158, 148, 17, 150, 564, 350, 307, 104, 550, 161, 27, 212, 94, 588, 358, 234, 187, 164, 274, 289, 147, 51, 97, 416, 134, 499, 336]. Analyzing [46, 440]. ancillary [207]. anisotropic [366]. Annane [282]. applicable [89]. Application [161, 337, 308, 241, 354, 314, 268, 133, 27, 535, 123, 612, 221, 305, 84, 336]. Applications [541, 422, 494, 523, 530, 566, 169, 414, 565]. applied [579]. applying [555, 404]. approach [487, 431, 476, 48, 162, 277, 391, 202, 330, 452, 536, 442, 499, 609]. approaches [602, 315]. approximants [543]. Approximate [188, 482]. Approximating [528]. approximation [519, 436]. approximations [422, 102, 31, 143]. Arabidopsis [46]. area [179, 587]. ARIMA [250]. arising [533, 535, 411]. arrays [397, 4]. arterial [26]. assessment [431, 24]. assessments [6]. assimilation [472]. assisted [472]. associated [291, 467, 345, 210]. association [573, 319, 308, 609]. assuming [550]. Astronomical [351, 163, 353]. astrophysical [352]. astrostatistics [349]. asymmetric [339]. asymmetry [233]. Asymptotic [324, 494, 185, 92, 100, 225]. Asymptotically [184, 427]. Asymptotics [24, 330]. Asynchronous [301]. atomic [106]. Attaching [369]. attraction [574]. attribute [136]. attribute-contaminated [136]. augmentation [266]. Author [12, 37, 85, 286, 72]. autocorrelated [363]. autocorrelation [216, 114]. Autoregressive [579, 179, 603, 185, 338, 273, 206]. auxiliary [448, 514]. average [343]. averaging [99].

B [282]. Background [356]. bands [148, 54]. Bandwidth [103, 220]. Bandwidth-based [103]. bank [269]. Barnard [607]. baryon [357]. based [554, 152, 374, 108, 505, 388, 604, 599, 538, 417, 13, 32, 461, 506, 573, 115, 454, 557, 103, 6, 469, 222, 588, 202, 78, 193, 4, 412, 92, 441, 316, 123, 449, 112, 206, 214, 551, 424, 339, 83, 221, 365, 578, 434, 170, 587, 609, 111]. baseline [559]. basis [76]. bathtub [234, 578]. bathtub-shaped [578]. Bayes [238, 294]. Bayesian [502, 614, 431, 476, 26, 306, 150, 480, 483, 307, 256, 577, 550, 503, 246, 197, 266, 226, 81, 358, 420, 559, 353, 370, 435, 264, 361, 583, 367, 84]. be [607, 563, 444, 295, 376]. behavior [315]. benchmark [148]. Bennett [377]. Berger [399]. Bering [467]. Bernoulli [53, 131]. Bernstein [312, 528]. best [402, 478, 395]. beta [186, 210]. beta- [210]. beta-type [186]. better [577, 402, 88]. between [192, 255, 520, 188, 407]. between-individual [520]. beyond [591]. Bhattacharyya [30]. Bias [434, 420, 43]. biased [551]. biases [260]. biasness [506]. bifurcating [185]. bilinear [140]. binary [440, 389, 105, 255, 171, 9, 141, 316, 558, 526]. binding [45]. binomial [36, 437, 579, 436, 610, 512, 542, 435, 119, 243]. Bioinformatics [39]. biological [41]. biorthogonal [165]. Birnbaum [516]. bispectrum [356]. Bivariate [459, 11, 26, 331, 454, 27, 516, 272, 562, 408, 142, 412, 89, 535, 214, 143]. Blazek [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 65]. block [421, 110, 584]. board [107, 38, 49, 55, 73, 86, 96, 116, 129, 137, 146, 153, 166, 173, 181, 189, 199, 208, 218, 228, 236, 242, 252, 270, 275, 287, 299, 309, 322, 327, 335, 342, 348, 372, 387, 394, 400, 405, 419, 425, 432, 439, 445, 453, 462, 474, 484, 489, 500, 509, 549, 515, 524, 539, 561, 568, 569, 580, 596, 615]. Book [341]. Bootstrap [495, 422, 110]. bound [577, 119, 243]. boundary [28]. bounds [494]. boxplot [233]. branch [269]. Bridges [255]. built [444]. Burr [175, 196, 380].

C [282]. calculation [15, 78]. can [295, 376]. cancer [47]. capability [315]. capture [465, 466]. capture-recapture-recovery [465, 466]. carbon [190]. carbon-resin [190]. cardioid [404]. care [597]. Carlo [46, 461, 364, 381]. case [314, 315, 518, 180, 111, 139]. Case-1 [139]. case-control [111]. cases [19]. catch [95]. catch-effort [95]. catchabilities [120]. categorical [447, 555, 134]. categories [98, 295, 377, 376]. causal [458]. causality [273]. Censored [126, 374, 177, 476, 613, 417, 18, 593, 411, 339, 97, 176, 293]. censoring [175, 517, 481, 567, 204, 375, 380, 139, 51, 176, 232, 170, 401]. Center [254]. certain [89]. Chain [46, 42, 461, 222, 364, 424, 381]. chains [381]. Change [603, 63, 443, 69, 66, 61, 68, 71, 67, 531, 64, 62, 70, 60, 65]. change-point [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 60, 65]. Chao [512]. chaotic [124]. Characterization [239, 454]. Characterizations [428]. characterizing [144]. chart [518]. check [525]. chemogenomic [308]. Chen [282]. Chernoff [8]. chi [427]. chi-square [427]. chloride [190]. choice [99, 5]. choices [555]. circular [31, 204]. claims [17]. class [115, 272, 588, 446]. classes [101, 334, 536]. Classification [59, 147, 230, 162, 43]. classifier [316]. clinical [195, 264]. clinico [150]. clinico-genomic [150]. clocks [188]. closeness [340, 401]. closure [532]. Cluster [554, 267, 172, 583, 83, 224]. Cluster-based [554]. clustered [440]. clustering [48, 162]. clusters [90]. CMB [156, 158]. coarsened [368]. Cocke [282]. codes [76]. coefficient [183, 592, 219, 589, 112, 194, 430, 521]. coefficients [216, 581]. Cohen [219, 295, 343, 321, 384, 407]. cointegrated [545]. cointegrating [545]. cointegration [545]. collapsible [558]. Combination [566, 235, 351, 514]. combining [386, 295, 377, 376]. Comments [433, 57]. common [563]. Comparative [203, 25]. compare [607]. Comparing [57, 246]. Comparison [373, 602, 233, 361, 16]. comparisons [192]. competing [564]. complementary [523]. complete [421, 454]. completeness [498]. Complexity [389]. compliance [458, 141]. component [158, 170]. components [183]. Composition [46, 328]. Computational [136]. computer [363]. Computing [7]. concentration [361]. concerning [506, 433]. Concomitants [130, 415, 331, 535]. condition [397, 240]. Conditional [423, 407, 429, 579, 179, 320, 454, 180, 318, 239, 257]. Conditionally [371]. conditioning [168]. conditions [207]. Confidence [385, 513, 581, 148, 54, 510, 188, 448, 508, 334, 316, 452, 399, 225, 346, 499]. confidence-set [452]. conforming [518]. Congress [259]. Congressional [259]. conjugate [258]. considerations [363]. Consistency [492, 560, 591]. consistent [263]. constant [536]. Constrained [378, 507]. constraint [523]. constraints [94]. Constructing [607, 238]. consumption [473]. contain [5]. contaminated [136]. contamination [541]. Contents [40, 155]. Context [45]. contingency [573, 319, 257]. continuous [403, 122, 614, 479, 528, 272, 239, 224]. continuous-discrete [479]. contrast [221]. contrast-based [221]. contributions [491]. control [601, 518, 111]. controlled [288]. controlling [522]. controls [402]. convergence [604, 397]. convex [607, 365]. Conway [502]. Copula [531, 443, 388, 449]. Copulas [11]. Cornish [426]. corona [161]. Corrected [92, 434]. correction [28, 231]. correlated [406, 440, 570, 496, 592, 323]. Correlation [562, 133, 357]. Correlational [260]. correlations [496, 611]. Corrigendum [393]. Cosmic [356]. count [570, 109, 442, 468]. counts [579]. covariance [492, 480, 211, 324, 370, 35]. Covariate [91]. covariates [534, 224]. coverage [577]. credible [577]. Credit [108]. criteria [32]. criterion [423, 588, 401]. critical [607]. Cross [32, 24]. cross-validated [24]. crossing [508, 305]. crossover [584]. cubic [30]. cumulant [251]. cumulants [237]. cumulative [582, 459]. Cure [139, 437]. Cure-rate [139]. curve [337, 520, 587]. curved [207]. curves [247, 391].

D [160]. DAG [538]. daily [361]. Data [52, 156, 406, 440, 570, 230, 157, 300, 374, 159, 352, 177, 290, 48, 613, 473, 150, 179, 209, 417, 18, 47, 362, 564, 191, 594, 350, 104, 409, 31, 413, 255, 470, 267, 246, 149, 523, 363, 41, 27, 163, 204, 180, 171, 109, 266, 588, 472, 14, 141, 172, 584, 20, 80, 135, 182, 234, 441, 316, 123, 35, 325, 393, 571, 551, 276, 339, 83, 265, 224, 97, 176, 536, 293, 416, 91, 308, 111, 360, 468]. datasets [447]. Dealing [264]. death [360]. debugging [606]. decomposed [50]. decomposition [441]. Deconvolution [350, 352]. decreased [295, 376]. decreasing [477]. deepest [493]. deformable [29]. delivery [6]. Demétrio [109]. Denoising [352, 452]. densities [237]. density [614, 302, 599, 479, 312, 28, 551, 51]. dependence [501, 492]. dependent [390, 135, 444, 485, 411, 45]. depth [106]. deriving [188, 136]. Description [36]. Design [149, 389, 429, 421, 483, 520, 584, 365, 240, 468]. designs [245, 363, 171, 212, 4, 373, 364, 583, 138, 288]. destructive [437]. detecting [233]. Detection [63, 69, 66, 61, 68, 71, 67, 64, 62, 164, 70, 60, 65, 157, 352, 44, 211, 603, 357, 98]. determinant [211]. determinants [261]. determination [192, 172, 401]. deterministic [461, 255, 369, 330]. development [190]. developments [390]. deviation [594]. diabetes [123]. Diagnosing [56, 26, 120]. Diagnostic [525, 502, 391]. diagnostics [604, 455, 182, 292]. dichotomous [172]. difference [318, 316]. differences [510, 585, 260]. Different [526, 244, 563, 214, 536]. differential [44]. Differentiated [543]. diffusions [598]. digraph [599]. dimension [540]. dimensional [230, 460, 492, 302, 493, 479, 608, 288]. Directional [416, 601]. Directionally [558]. directions [542]. Dirichlet [337, 21, 559]. Discovering [19]. discovery [410, 601, 45]. discrepancy [32]. Discrete [606, 345, 196, 403, 457, 479, 248, 259, 548, 553, 536]. Discriminant [187, 487]. Discussion [69, 66, 283, 281, 68, 280, 282, 285, 71, 67, 284, 70, 65, 63, 61, 286, 64, 62]. disease [277, 368]. disparity [182]. dispersed [570, 579, 100]. dispersion [398]. distance [140, 193, 111]. distinguishing [47]. distortion [552]. distributed [301]. Distribution [541, 175, 244, 58, 127, 203, 450, 457, 502, 374, 331, 544, 481, 593, 297, 329, 102, 409, 113, 550, 454, 557, 385, 523, 504, 21, 235, 248, 186, 610, 94, 548, 344, 562, 15, 451, 92, 118, 334, 535, 375, 441, 565, 380, 433, 100, 174, 612, 424, 83, 221, 578, 434, 475, 572, 416, 426, 427, 499, 336]. Distribution-free [541, 572]. distributional [417]. distributions [144, 58, 403, 291, 130, 476, 493, 533, 209, 417, 314, 34, 313, 605, 229, 537, 115, 567, 235, 101, 332, 516, 196, 272, 226, 136, 89, 446, 513, 239, 590, 201, 79, 558, 468, 210]. divergence [576, 506, 573, 338]. document [301]. Dodge [390]. Does [259]. domain [182, 330, 165]. domains [574]. Donoho [494]. dose [195]. Double [557, 220, 518]. Doubly [14, 125, 571]. down [311]. drift [365, 598]. dropout [526]. drug [6]. dry [205]. duck [468]. due [526]. Dynamic [319, 550, 160, 362, 503, 315, 268, 603, 582, 459, 124]. dynamic-behavior [315]. dynamical [472]. dynamics [487, 187].

eastern [467]. Ecological [262, 472]. ecology [542]. Edge [599]. Edgeworth [383]. editor [253]. Editorial [107, 463, 1, 359, 349, 154, 38, 49, 55, 73, 86, 96, 116, 129, 137, 146, 153, 166, 173, 181, 189, 199, 208, 218, 228, 236, 242, 252, 270, 275, 287, 299, 309, 322, 327, 335, 342, 348, 372, 387, 394, 400, 405, 419, 425, 432, 439, 445, 453, 462, 474, 484, 489, 500, 509, 549, 515, 524, 539, 561, 568, 569, 580, 596, 615]. effect [458, 377]. Effectiveness [274]. effects [328, 120, 555, 83, 75]. efficiencies [278]. efficiency [310, 57]. Efficient [456, 297, 329, 592, 171, 4, 499, 184, 142]. effort [95]. eigenvalues [100]. element [82]. elementary [8]. ellipsoids [581]. elliptical [34, 455, 235, 333, 590]. emission [91]. Empirical [487, 143, 422, 505, 448, 441, 294, 126, 426, 521]. Emulator [472]. Emulator-assisted [472]. endogenous [268]. endpoints [409]. enhance [447]. ensemble [162]. Entropy [389, 245, 582, 588, 511, 459, 600, 111]. entropy-based [588]. epilogue [254]. episodes [597]. epoch [157]. equality [421, 216]. equation [550, 215]. equicorrelated [455]. equivalence [15]. Equivalences [382]. ergodic [613]. Erratum [298]. Error [41, 87, 522, 550, 534, 226, 215, 112, 119, 243]. error-prone [534]. errors [498, 552, 455, 592, 531, 215, 273, 589]. errors-in-variables [215]. estimates [458, 496]. Estimating [18, 223, 546, 133, 610, 23, 95, 380, 124, 368, 249, 501, 41, 80]. Estimation [175, 203, 450, 374, 198, 602, 331, 593, 523, 145, 222, 141, 195, 250, 395, 289, 578, 244, 10, 301, 410, 354, 271, 388, 302, 303, 613, 479, 456, 13, 297, 329, 24, 480, 313, 504, 603, 140, 379, 248, 592, 28, 74, 584, 608, 378, 535, 433, 139, 551, 51, 176, 213, 132, 293, 305, 84, 587, 323, 468]. Estimations [21, 540]. estimator [24, 211, 125, 330, 356, 294, 206, 22]. Estimators [449, 168, 184, 544, 310, 530, 529, 312, 520, 357, 560, 512, 392, 475]. EUV [161]. evaluating [315]. Evaluation [283, 286, 280, 282, 281, 298, 279, 285, 284]. events [290, 317, 392, 323]. Evolutionary [364, 483]. Exact [517, 567, 235, 417, 350, 438]. example [302, 399]. Excess [117, 442]. exchange [50]. exchangeable [224]. existence [207]. exoplanet [358]. expansion [100]. Expansions [427, 494, 383, 426]. experimental [363]. experiments [190, 192, 4, 5]. experts [267]. Explicit [438]. exploration [495]. exponent [124]. exponential [291, 331, 517, 417, 481, 593, 567, 557, 385, 523, 30, 586, 510, 266, 328, 207, 344, 562, 451, 334, 132]. exponential-family [328]. exponentiality [88]. exponentiated [244, 523, 375]. exposure [6]. expression [42, 48, 44, 41, 438]. Extended [58, 36, 307, 520, 535, 289]. Extending [518]. extension [344, 459]. Extensions [390, 345, 143]. exterior [7]. external [365]. extraction [91]. extremal [501, 135]. Extreme [249, 102, 409].

factor [45]. failure [556, 564, 234, 536]. Fallacious [399]. fallible [316]. false [410, 601]. families [291, 30, 586, 210]. family [599, 573, 537, 101, 328, 207, 435, 376]. familywise [522]. Farlie [130, 535, 224].

Fast [165, 356]. Fay [423]. Feature [47]. fibration [355]. fidelity [350]. fields [371]. Filtered [547]. final [112]. Finite [410, 58, 420, 482]. First [102]. fish [541]. Fisher [375, 482, 426]. fit [602, 229, 33, 20, 98, 89]. Fitting [247, 468, 159]. Fixed [330, 494, 508, 294, 225]. Fixed-domain [330]. fixed-size [508]. Florida [283, 281, 286, 298, 280, 282, 285, 284, 279]. flow [161]. fMRI [192]. follow [317]. follow-up [317]. following [213]. food [473]. Forecasting [151]. Foreword [2]. form [438]. formula [168]. Forward [608, 160]. foundations [497]. fractional [250, 547]. frailty [314]. framework [466]. free [541, 572]. frequencies [257]. frequency [584, 147]. frequentist [577]. Fully [307, 431]. function [36, 479, 454, 125, 379, 180, 357, 446, 392, 176]. functional [230, 303]. functionals [426]. functions [144, 505, 519, 469, 528, 318, 435]. fundamental [147]. Further [532]. fusion [163, 353].

G [131, 282]. Gaia [354]. galaxy [357]. gamma [374, 478, 210, 336]. gamma-generated [210]. GARCH [10, 531, 511]. Gaussian [105, 13, 494, 261, 480, 229, 330, 598, 609]. gene [42, 48, 44, 41]. General [505, 497, 564, 498, 320, 140, 204, 101, 333, 202, 292, 565, 339, 416]. generalised [466]. generalizations [446]. Generalized [570, 481, 144, 203, 450, 389, 130, 428, 158, 26, 533, 517, 522, 593, 113, 550, 567, 385, 529, 27, 312, 344, 188, 373, 591, 441, 174, 612, 57, 210, 336]. generated [210]. genes [43]. genetic [50]. genomic [150]. geocoding [368]. geometrical [152]. geometry [251]. Gini [585]. given [293]. glitches [164]. Going [591]. gold [316]. Goldstein [377]. goodness [602, 20, 98, 89]. goodness-of-fit [602, 89]. gradient [398].

Granger [273]. graph [599, 266, 328]. graphical [42, 261]. graphs [320]. grid [354]. Group [212, 464, 192]. growth [290, 247, 261, 337, 520]. guarantee [607]. Guest [253, 154]. Gulati [282]. Gumbel [130, 535, 224].

Haar [165]. half [557]. Hamid

[283, 281, 280, 282, 285, 284]. handling [526]. harmonics [355]. having [478]. Hazard [379, 176, 553]. health [179]. HeartCast [597]. heavy [476, 17, 525]. heavy-tailed [476, 17]. Hellinger [140]. Hermite [209]. Herriot [423]. heterogeneity [464, 75]. heteroscedastic [184, 215]. heteroscedasticity [510]. Hidden [160, 614, 223]. Hierarchical [114, 307, 467]. High [302, 518, 608]. High-dimensional [302]. Higher [237, 251, 321, 384]. Higher-order [237]. Hinde [109]. HNBRUE [241]. Hoeffding [143]. Hölder [164]. Holm [311]. Homogeneity [586, 325, 393]. homogeneous [371, 14, 176, 381]. Hopf [355]. human [223]. Hurricane [283, 286, 280, 282, 302, 281, 298, 279, 285, 14, 284]. hybrid [517, 417, 481, 567]. hyperspectral [352, 353]. hypotensive [597]. hypotheses [576, 598]. Hypothesis [192, 241, 165].

ice [467]. ice-associated [467]. Idea [238]. ideas [365]. identical [411]. identifiability [458]. Identification [320]. identifiers [497]. II [172, 4, 375, 380, 411, 170]. Illustration [409]. Image [29, 238, 351]. images [163]. immunes [108]. implementation [226, 35]. Implications [468]. importance [3]. important [74]. Improved [475, 334, 514]. improving [132]. imputation [93, 555, 265]. incomplete [588, 584, 526, 368]. Incorporating [514, 268, 41]. increased [295, 376]. independence [505, 320, 143]. Independent [485, 59, 226, 444, 356].

Index [12, 37, 85, 534, 608, 135]. indexed [505]. indicator [51]. indicators [74, 176]. individual [520, 120]. induced [527]. industrial [180]. inequalities [241, 79, 407]. infant [147, 360]. Inference [193, 221, 265, 611, 406, 502, 614, 447, 604, 26, 517, 386, 480, 256, 552, 115, 503, 262, 567, 246, 103, 185, 420, 559, 589, 224, 210]. Inferences [93, 332]. inferiority [15]. inferring [42]. infinite [493]. infinitely [100]. inflated [570, 502, 118, 565]. inflation [332]. Influence [292, 502, 455, 20]. influential [19, 98]. information [63, 69, 66, 61, 423, 93, 68, 71, 67, 251, 448, 64, 62, 70, 375, 60, 65, 514]. Informative [263, 323]. inhomogeneous [483]. initial [223, 171]. innovations [318, 250]. instrumental [430]. integer [575, 610]. integrability [397]. intensity [125]. intensive [363, 597]. interaction [555]. interactions [304]. interdependence [268]. intermediate [574]. intermittent [104]. interpolation [209]. interpolations [369]. Interpreting [362]. Interrelated [48]. interval [556, 496, 436, 202, 139, 399]. intervals [577, 385, 510, 188, 448, 541, 334, 316]. Introduction [39, 300, 253]. intrusions [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 60, 65]. invariance [537]. Invariant [34]. Inverse [232, 429, 431, 303, 229, 248, 612]. inverted [481, 593]. irregularly [159, 191]. isotones [167]. Issue [40, 359, 39, 300, 349, 253, 490]. iterated [77].

J [282]. J.-P [282]. Jacobians [7]. Jeffreys [291]. Jin [494]. Jolly [464].

kappa [219, 194, 295, 343, 321, 384, 407]. kappas [382, 376, 407]. Kendall [449]. kernel [149, 28, 33, 51, 587, 514, 609]. kernel-based [609]. kernels [345]. Kesar [490, 491]. Kibria [282]. Kim [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 65]. knowledge [41]. known [128]. Kronecker [35]. Kullback [32, 506]. Kumaraswamy [186, 336].

L [293]. L-moment [293]. lack [33].

language [147]. Laplace [457, 174]. Large [604, 492, 397, 353, 305]. large-margin [305]. latency [260]. Latent [588, 602, 437, 3, 480, 267, 94]. lattice [367]. lattices [440]. Law [77, 290, 92]. laws [574, 397, 193]. learning [127, 306, 47]. least [91]. left [97]. Leibler [32, 506]. Length [36]. lengths [110]. lesson [127]. level [594, 212, 555]. levels [195, 361]. Leverage [278]. life [175, 18, 477, 557, 523, 101, 566]. lifetime [578]. Likelihood [13, 316, 224, 487, 443, 604, 517, 544, 567, 436, 466, 520, 81, 448, 330, 206, 276, 126, 132, 442, 521]. Likelihood-based [13, 316]. limit [574, 77]. Linear [94, 389, 486, 476, 26, 32, 19, 258, 256, 498, 552, 324, 54, 529, 534, 27, 488, 235, 371, 33, 227, 446, 219, 373, 591, 395, 540, 525, 521]. linearity [227]. linearly [194, 384]. Link [223, 435]. Link-tracing [223]. list [40, 120]. Local [185, 398, 240, 576, 33, 591]. Location [590, 121, 310, 510, 142, 412, 399, 75]. Location-scale [590]. locations [368]. Loci [46]. Log [289, 258, 256, 22]. log-linear [258, 256]. log-normal [22]. Log-Weibull [289]. logarithm [77]. logarithmic [519]. logdensity [543]. Logistic [604, 203, 450, 535, 123, 111]. long [250, 84]. long-memory [250]. longitudinal [594, 104, 222, 526]. Looking [590]. Loss [283, 286, 280, 282, 464, 281, 298, 279, 285, 284, 339]. losses [283, 281, 286, 298, 279, 280, 282, 285, 284]. lower [577]. LPQE [532]. LQE [595]. LR [398]. Lyapunov [124].

M [282, 131]. machine [306, 47, 307]. machines [514]. macro [538, 19]. macro-DAG [538]. macroeconomic [261]. macrolevel [114]. MAOSA [44]. mapping [202]. margin [305]. marginal [424]. Markov [614, 46, 160, 461, 546, 222, 74, 123, 364, 424, 381, 442]. Markov-modulated [546]. Marshall [562]. martingale [318]. masking [497]. mass [47]. matched [214, 325, 393]. matched-pairs [325, 393]. matching [429, 78]. matrices [183, 324, 30, 520]. Matrix [79, 492, 375]. Maximum [520, 511, 443, 494, 330, 132]. Maxwell [502]. may [563, 444]. MCMC [483, 226]. MDL [36]. mean [494, 18, 397, 566, 610, 318, 446, 508, 395, 585, 260, 22]. means [418]. measure [506, 106]. measurement [87, 552, 226, 273, 589]. measurements [59, 141, 572, 521]. measures [501, 573, 277, 123, 35, 276, 111]. Median [97, 340, 494, 296, 563, 126, 232]. mellitus [123]. Memorial [490]. memory [250, 84]. mercury [541]. meta [499]. meta-analysis [499]. meteorological [27]. method [388, 332, 345, 212, 482, 471]. Methodol [298, 393]. methodologies [508]. methodology [491, 169]. methods [63, 159, 233, 69, 247, 66, 61, 363, 68, 71, 67, 110, 358, 95, 64, 200, 62, 70, 253, 526, 60, 351, 65]. Microwave [356]. minimal [460, 240]. Minimum [36, 211]. mining [300, 308]. MINITAB [19]. misclassified [20]. Mises [113, 441]. missing [456, 564, 104, 93, 266, 264, 526, 265, 51, 176]. misspecification [373]. Mixed [205, 594, 212, 390, 82]. mixed-level [212]. mixing [418]. Mixture [3, 94, 58, 614, 410, 538, 267, 337, 235, 516, 512, 590, 132]. mixtures [313, 605, 482, 265]. MLEs [434]. Möbius [404]. mode [258, 537, 454, 441]. mode-invariance [537]. Model [46, 99, 286, 441, 42, 440, 418, 614, 431, 354, 556, 437, 538, 613, 579, 473, 456, 494, 32, 314, 34, 498, 409, 324, 267, 298, 279, 423, 268, 54, 145, 467, 534, 21, 41, 222, 277, 27, 259, 285, 610, 520, 226, 202, 114, 120, 512, 80, 182, 451, 234, 373, 591, 118, 215, 292, 356, 589, 527, 289, 339, 367, 399, 224, 51, 351, 82, 346, 464]. model-independent [356]. Modeling [261, 50, 570, 502, 108, 160, 606, 319, 470, 6, 590, 360]. Modelling [191, 180, 302]. Models [152, 283, 280, 282, 10, 301, 389, 410, 87, 443, 486, 388, 184, 148, 602, 3, 476, 26, 150, 362, 205, 594, 258, 261, 480, 314, 605, 256, 552, 320, 31, 255, 503, 281, 396, 603, 140, 529, 337, 465, 455, 197, 592, 466, 109, 266, 94, 328, 318, 338, 531, 511, 333, 398, 371, 33, 227, 14, 608, 560, 408, 373, 438, 151, 378, 284, 273, 56, 330, 370, 449, 581, 435, 361, 571, 424, 367, 45, 132, 293, 540, 84, 442, 111, 430, 521]. modifications [167]. Modified [507, 430, 54, 202]. Modifying [220]. modulated [546]. moisture [470]. Moment [241, 102, 293]. moments [237, 388, 544, 530, 332, 528]. monitoring [556, 364, 365, 323]. monotone [308]. Monte [46, 461, 364, 381]. Morgenstern [130, 331, 535, 224]. morphological [158]. MOSS [258]. most [460]. movement [469]. movement-based [469]. movie [161]. Moving [249]. MRFs [13]. Multi [163, 230, 157, 594, 413, 555, 376]. multi-dimensional [230]. multi-epoch [157]. multi-level [594, 555]. multi-rater [376]. multi-sample [413]. Multi-source [163]. multicollinearity [488]. Multidimensional [135]. multidimensions [19]. Multilevel [570]. multinomial [256, 92]. multinomials [482]. multiparameter [586]. Multiple [496, 508, 59, 159, 53, 93, 145, 488, 545, 391, 78, 555, 265, 382, 16]. multiple-threshold [145]. Multivariate [333, 553, 551, 198, 310, 34, 550, 315, 455, 226, 472, 508, 201, 378, 8, 292, 79, 449, 35, 558, 276, 225, 383]. MUSE [352].

natural [291]. Navigating [555]. NBU [101]. nearest [288]. necessary [240]. negative [437, 579, 542]. nested [314, 138]. network [473, 470, 267, 268, 328, 364, 365, 499]. networks [42, 41, 50, 269]. neural [50]. NGINAR [237]. noise [350, 124, 165]. nomination [541]. Non [614, 225, 305, 458, 47, 125, 518, 371, 14, 15, 141, 411]. Non-asymptotic [225]. non-cancer [47]. non-compliance [458, 141]. non-conforming [518]. Non-crossing [305]. non-homogeneous [371, 14]. non-identical [411]. non-inferiority [15]. Non-parametric [614, 125]. noncompliance [195, 264]. Nonconvex [540]. Nonlinear [605, 320, 531, 472, 33, 560, 589, 225]. Nonparametric [613, 479, 105, 74, 370, 392, 121, 178, 198, 220, 184, 247, 483, 413, 503, 603, 103, 379, 110, 391, 142, 132]. nonserial [8]. normal [271, 313, 605, 229, 550, 610, 508, 559, 292, 22, 416]. normal/independent [226]. normality [185, 167]. normalization [574]. normals [337, 265]. note [574, 415, 458, 131, 312, 510, 219, 347, 194, 213]. Notes [584]. notion [477]. novel [609]. numbers [397].

Object [157]. observables [339]. observations [358, 353, 97]. observed [191, 466]. obstruction [26]. odd [340, 544, 296]. odds [249, 429, 550, 111]. offset [188]. Olkin [562]. one [454, 347]. optical [161]. Optimal [353, 5, 170, 230, 483, 212, 110, 364, 583, 288, 240, 401]. optimality [311, 240]. option [5]. oracle [591]. order [144, 237, 340, 574, 130, 428, 331, 102, 235, 532, 566, 117, 251, 535, 411, 611]. Ordered [386, 209, 4]. ordering [582, 595, 600]. orderings [585]. ordinal [194]. oriented [258]. Orthogonal [190, 520, 4]. Oscillation [161, 357]. other [454, 235]. outcome [526]. outcomes [264]. outlier [488, 135, 497]. outlier-proneness [135]. Outliers [211, 290, 233, 329, 433]. over-dispersed [579]. overdispersed [109, 114]. ozone [361].

P [282, 145, 266]. p-star [266]. pair [202, 214]. paired [177, 392]. pairs [429, 325, 393]. Pakistan [365]. panel [362]. papers [433]. parallel [170]. parallelism [572]. Parameter [354, 465, 323, 374, 331, 517, 577, 21, 332, 531, 451, 118, 334, 375, 565, 578]. parameterization [333, 292]. parameterizations [558]. parameters [203, 290, 546, 248, 510, 23, 448, 478, 250, 294, 380, 433, 225, 475, 346, 111]. parametric [614, 501, 125, 435, 293]. Pareto [244, 297, 329, 504, 548, 196, 23, 433, 475]. Partial [454, 552, 132, 426, 91, 442]. Partially [476, 175, 178, 486, 534, 466, 227, 589, 325, 393, 540, 521]. partisan [259]. partitioned [3]. path [308]. pattern [564, 91]. patterned [178]. patterns [147]. Peeking [19]. Penalized [230, 540]. penalty [36, 529, 507]. percentile [477]. percutaneous [6]. Performance [294, 24]. perinatal [264]. periodic [606, 125, 14]. periods [11]. permutation [25, 200]. perspective [386, 52]. phase [172]. phases [53]. Pinelli [282]. Pitman [401]. pivotal [188]. plan [557]. plans [390, 170]. PLS [91]. plug [110]. plug-in [110]. plumes [160]. point [63, 443, 493, 69, 66, 61, 68, 71, 67, 357, 80, 64, 62, 70, 60, 471, 65]. points [77]. Poisson [127, 570, 502, 350, 256, 246, 546, 125, 114, 14, 15, 584, 565, 547, 119, 243, 471, 82, 346, 165]. Poisson-binomial [119, 243]. polar [160]. policy [53, 131]. political [268]. polynomial [312]. polynomials [209, 528]. population [340, 168, 223, 23, 95, 420, 100, 611]. populations [421, 478]. position [267]. positive [410, 193]. positron [91]. Posterior [435, 221]. Powell [282]. power [574, 192, 290, 557, 398, 118, 609, 75]. power-based [609]. power-law [290]. powerful [460]. Practical [363]. practice [211]. precise [22]. precision [271]. preconditioning [305]. Predicting [283, 281, 286, 298, 279, 280, 282, 285, 597, 284, 269, 339]. Prediction [277, 593, 112, 514]. predictive [420, 559]. predictor [395]. predictors [16]. Preface [490]. preference [5]. prehistory [254]. prequel [254]. presence [329, 488, 555, 433]. preserve [257]. pretest [529]. Prevention [264]. primary [581]. Principal [183]. Prior [389, 105, 179, 307, 132]. priors [258, 350]. Probabilistic [6, 231, 255]. probabilities [340, 17, 396]. probability [205, 577, 245, 169, 305]. probit [87, 408]. problem [121, 413, 563, 142]. problems [431, 303, 412, 542, 507]. procedure [178, 44, 311]. procedures [136, 611]. Process [470, 237, 575, 105, 480, 546, 559, 330]. processes [487, 422, 505, 315, 366, 185, 518, 472, 74, 187, 547, 206, 471]. Processing [352, 560]. product [35]. products [7]. Professor [490]. profile [126]. profiles [42]. profitability [269]. progressive [517, 567, 380, 339, 401]. progressively [374, 417, 593, 411]. projected [416]. prone [534]. proneness [135]. proof [8]. Properties [313, 565, 21, 532, 398, 239, 411, 527]. property [311, 296, 591]. proportion [436, 518, 141, 420]. proportional [288]. proportions [607, 36]. propriety [435]. proximity [182]. Public [283, 281, 286, 280, 282, 284, 298, 279, 285].

Quadratic [303, 193]. quadratically [384]. qualitative [304]. quality [518, 52]. quantal [148]. quantile [422, 582, 563, 448, 446, 600]. Quantitative [46]. quantities [251, 188]. Quasi [81, 240]. quasi-necessary [240]. queue [131]. queueing [53].

rainfall [365]. random

[599, 131, 77, 455, 235, 528, 266, 328, 371, 136, 169, 444, 411, 526, 83, 404]. random-effects [83]. randomised [421]. randomization [172]. randomized [458, 141, 195, 264]. randomly [397]. Rank [138, 587, 386, 133, 545, 472, 412]. Rank-based [587]. Ranked [249, 413, 214]. ranked-set [413]. ranking [214]. Rasch [120]. Rate [379, 556, 437, 601, 234, 553, 139, 536]. rater [376]. raters [382]. rates [410, 522, 246, 50]. ratio [429, 271, 443, 78, 141, 195, 169, 128, 276, 111]. Ratios [444, 550, 485]. ray [162]. re [554]. re-weighted [554]. reader [391]. realign [259]. recapture [465, 466]. recognition [178, 29]. reconstruction [158, 164]. record [578]. records [415, 434]. recovery [465, 466]. recurrence [406]. recurrent [290, 392, 323]. reduced [472]. reduced-rank [472]. reduction [533]. redundancy [465]. regimes [259]. region [508]. regions [607, 225, 346]. Regression [564, 144, 554, 570, 502, 428, 198, 220, 486, 604, 184, 476, 613, 519, 105, 456, 32, 19, 247, 483, 605, 54, 396, 455, 222, 488, 278, 109, 33, 202, 9, 448, 292, 123, 581, 435, 289, 126, 97, 232, 225, 240, 91, 442, 111]. regressive [277]. regularization [471]. regularly [191]. regulatory [42]. related [185, 610]. relationships [545]. relative [133, 119, 243]. relevance [307]. reliability [450, 606, 247, 566, 74, 234, 578]. removal [190, 165]. renal [26]. Rényi [216, 600]. repeated [277, 141, 123, 35, 276, 572, 521]. replacement [90]. replications [240]. representation [152]. representations [21]. Resampling [200]. residential [283, 281, 286, 298, 279, 280, 282, 285, 284]. Residual [83, 477, 21, 582, 566, 446, 459, 289].

Residual-based [83]. resin [190]. resolution [163]. resource [469]. response [122, 148, 171, 9, 260, 72]. responses [389, 456, 286]. restoration [351]. restricted [46, 131, 577, 475]. restrictions [611]. result [530]. results [417, 532, 595, 20, 378, 8, 600]. return [11]. reversed [379]. Review [217, 326, 542, 513, 341]. Revisiting [399]. rich [127]. ridge [540]. ridit [200]. rigged [399]. right [613, 18, 97, 176]. risk [108, 24, 6, 78, 195, 316, 368]. risk-based [6]. risks [564, 523]. robust [310, 226, 590, 57, 305]. robustness [497]. ROC [391, 587]. ROC-curves [391]. root [197]. rounding [231]. Rozovskii [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 65]. ruin [17]. rule [122, 110]. RVS [354].

S [283, 281, 280, 282, 285, 284]. S.-C [282]. Saddlepoint [31]. Sample [15, 78, 172, 244, 340, 121, 492, 192, 604, 233, 494, 506, 223, 413, 296, 117, 563, 142, 347, 316, 395, 128, 572]. sampled [159]. samples [47, 518, 541, 43]. Sampling [249, 59, 429, 3, 90, 223, 245, 557, 21, 518, 136, 390, 214, 365, 288]. Saunders [516]. Savage [8]. SCA [302]. SCAD [507]. scalar [553]. scale [313, 605, 537, 413, 469, 563, 590, 334, 475]. scales [194]. scaling [9]. schedule [53, 131, 606]. scheduling [358]. scheme [59, 437, 533, 517, 204, 274, 401]. schemes [244, 567, 214]. Schilling [390]. science [268]. Sciences [254, 262, 253]. score [398, 438]. scoring [230, 482]. screening [212]. sea [468, 467]. seal [467]. search [258]. seasonal [546, 250, 56]. Seber [464]. Seber-Tag-Loss [464]. SEE [430]. seemingly [581]. selected [23, 542]. Selecting [478, 110]. Selection [591, 43, 47, 32, 24, 314, 402, 304, 469, 488, 608, 420, 507, 441, 263, 367, 471, 213, 609, 430]. selector [220]. semi [306, 74]. semi-Markov [74]. semi-supervised [306]. Semiparametric [406, 115, 521, 443, 388, 408]. Sensitivity [17, 104, 391, 559, 289]. sensor [470]. separability [276]. separate [59, 75]. sepsis [264]. sequence [527]. Sequential [522, 542, 598, 178, 63, 69, 66, 223, 483, 61, 68, 71, 67, 64, 62, 169, 508, 70, 60, 225, 65]. serial [8]. serially [592]. series [152, 198, 575, 579, 140, 318, 338, 182, 438, 118, 56, 76, 134, 525, 442, 124]. Set [249, 131, 413, 452, 214]. set-up [131]. sets [607, 290, 5]. setting [519]. several [421, 378]. shape [478]. shaped [409, 578]. Sharp [494]. Sheppard [231]. shifts [56]. Shrinkage [529]. Shtarkov [291]. sib [202]. sib-pair [202]. sign [214]. signal [560, 164, 452]. signed [412]. significance [75]. similarity [147]. simple [168, 436, 54, 451, 536]. Simultaneous [178, 148, 402, 346, 340, 510, 318]. Singh [490, 491]. single [121, 534, 608]. single-index [534]. sinusoids [159]. sites [223, 45]. six [16]. size [192, 168, 233, 223, 328, 398, 15, 78, 172, 508, 551, 288]. size-biased [551]. sizes [340, 296, 95, 395, 224, 225, 82]. Skew [87, 313, 457, 410, 575, 34, 605, 537, 550, 235, 226, 590]. skew-elliptical [235, 590]. skew-mixture [410]. skew-normal [550, 226]. skew-normal/independent [226]. Skew-probit [87]. Slope [54]. slopes [412]. small [518]. smooth [594, 379]. Smoothed [448]. smoothers [149, 33, 16]. smoothing [220]. smoothly [259]. Social [254, 473, 262, 267, 253]. software [606, 247]. soil [470]. solar [161]. solution [591]. solutions [190]. Some [366, 595, 414, 527, 167, 600, 291, 533, 417, 31, 186, 74, 378]. source [163]. sources [352, 162, 80]. South [264]. Soweto [264]. space [362, 191, 503, 4, 438, 56]. space-time [362]. spaces [46, 493, 577, 251]. spacings [117]. sparsity [156, 350]. Spatial [365, 359, 369, 149, 268, 467, 371, 80, 370, 367, 221, 471, 91, 368]. spatial-temporal [91]. spatially [440, 179, 363]. Spatio [360, 472]. Spatio-temporal [360, 472]. Special [40, 253, 359, 39, 300, 193, 349, 274]. species [467]. specific [5]. specification [454, 83]. specificity [391]. specified [88]. spectra [354]. Spectral [182, 237, 501]. spectrometry [47]. spectrum [198]. spells [205]. sphere [366, 174]. spherical [355]. spherically [498]. Spin [355]. splines [227]. spoken [147]. square [112, 427]. squares [91]. stable [487, 193, 250, 187]. stacking [409]. stage [93]. standard [316]. star [266]. Stat [298, 393]. state [283, 503, 281, 286, 298, 279, 280, 282, 285, 438, 284, 56]. state-space [503]. stated [5]. stationarity [424]. stationary [613]. statistic [460, 25]. Statistical [486, 552, 589, 491, 300, 473, 310, 52, 212, 542, 253, 468]. Statistics [254, 144, 340, 359, 574, 130, 428, 331, 386, 102, 216, 235, 207, 92, 535, 411, 221, 138, 427]. status [277]. step [175, 556, 311, 523, 451]. step-down [311]. step-stress [175, 556, 523, 451]. Stochastic [197, 258, 527, 84]. strength [450, 578]. stress [175, 450, 556, 523, 451, 578]. stress-strength [450]. Strong [422, 397]. structural [324, 215]. structure [538, 307, 94, 562, 123, 35]. structures [501, 283, 281, 286, 298, 279, 280, 282, 285, 284]. Studies [251, 152]. study [203, 317, 544, 315, 180, 25]. studying [497]. subgroups [418]. subject [246]. subpopulation [308]. Subset [488, 278]. subsets [98]. successive [510]. sudden [360]. sufficient [460, 240]. sum [426]. sums [77, 397]. super [163]. super-resolution [163]. supersaturated [389, 212]. supervised [306, 43]. support [306]. sure [77]. survey [468]. surveys [357]. Survival [51, 108, 150, 293, 336]. survivor [392]. SVM [305]. swamping [497]. switching [259]. Symmetric [601, 498, 174]. symmetry [201]. symplectic [355]. syndrome [360]. synthesis [350]. Synthesizing [447]. Synthetic [257]. system [42, 53, 550, 274, 170]. Systematic [594]. systems [63, 69, 66, 61, 68, 71, 67, 545, 64, 62, 70, 60, 65]. SZ [158].

tables [573, 319, 257, 321, 384]. Tag [464]. tail [525]. tailed [476, 17]. tampered [556]. TAR [151]. Tartakovsky [63, 69, 66, 61, 68, 71, 67, 64, 62, 70, 65]. tau [449, 308]. tau-path [308]. techniques [51]. technology [264]. templates [29]. Temporal [469, 472, 91, 360]. terms [550]. Test [429, 460, 421, 443, 477, 413, 563, 318, 338, 531, 511, 33, 202, 227, 142, 412, 89, 169, 414, 347, 35, 214, 128, 276, 424, 138, 381, 170, 572, 143, 308, 609]. Testing [418, 290, 317, 545, 120, 201, 118, 585, 241, 576, 192, 496, 216, 534, 586, 15, 20, 98, 182, 88, 273, 76, 598, 536, 111, 165]. Tests [177, 229, 325, 393, 175, 607, 121, 178, 505, 486, 522, 506, 34, 557, 523, 222, 197, 398, 167, 521, 75]. thaliana} [46]. their [167]. theorems [397]. theoretical [602]. Theory [133, 169, 366]. there [16]. three [374, 467, 375]. three-parameter [374, 375]. threshold [145, 545]. Thurstone [9]. time [406, 198, 575, 579, 131, 606, 18, 362, 191, 477, 523, 140, 318, 338, 4, 182, 438, 414, 56, 76, 134, 525, 442, 124]. times [317]. tissue [43]. tolerance [541]. tolerant [41]. tomography [91]. top [43]. topic [301]. toric [381]. total [477, 414]. tracing [223]. tracks [302]. tractability [186]. traffic [89]. Trait [46, 602]. transcription [45]. transdermal [6]. transform [477]. Transformation [234, 537, 571, 404]. Transformations [383, 461]. transforms [414]. Treatment [200, 122, 263]. treatments [402]. tree [150]. Trees [160]. trial [195, 264]. trials [458, 391, 141, 172]. trichotomous [195]. tridiagonal [321, 384]. trimmed [449]. trinomial [612]. trivariate [533, 235]. trophic [473]. truncated [557, 548, 571]. Tukey [106]. Two [75, 607, 460, 122, 233, 48, 479, 53, 246, 93, 235, 563, 188, 142, 478, 451, 347, 334, 316, 565, 581, 128, 433, 257, 578, 288, 572, 143]. two-dimensional [460, 288]. two-parameter [451, 334, 565, 578]. two-sample [563, 142, 347, 316, 128, 572]. two-stage [93]. two-treatment [122]. two-way [48, 257]. type [175, 58, 331, 186, 4, 375, 380, 383, 417, 411, 170]. Type-I [417]. Type-II [411, 170]. types [599, 564].

UAV [523]. ultrahigh [540]. unbalanced [276, 138]. unbiased [395, 213]. uncertain [603]. Uncertainty [357, 261, 369]. uncorrelated [440]. underreporting [246]. undetected [317]. unequal [245, 478, 395]. unified [391, 8]. uniform [243]. Uniformly [460]. uniqueness [530, 591]. unit [197]. unit-root [197]. units [597, 4]. Univariate [315, 233, 151, 8]. Universal [76]. unknown [347]. unrelated [581]. unrestricted [46]. unspecified [308]. unsupervised [162]. upper [578]. use [10, 224]. used [88]. useful [378]. Using [33, 227, 111, 175, 127, 249, 7, 389, 178, 158, 46, 160, 476, 331, 105, 18, 247, 261, 350, 320, 413, 470, 161, 246, 363, 21, 528, 266, 4, 50, 234, 164, 56, 264, 214, 269, 585, 365, 265, 51, 351, 143, 514].

vacation [53, 131]. validated [24]. validation [32, 246]. validity [75]. value [102]. valued [575]. values [386, 104, 408, 578]. VAR [320, 10]. variability [585]. Variable [304, 471, 235, 507, 430]. variables [3, 235, 528, 136, 444, 215, 411, 365, 404]. variance [271, 168, 324, 180, 318, 239, 347, 79, 75]. variance-covariance [324]. variances [421, 128, 57]. variation [183, 469, 368]. variational [351]. varying [418, 592, 589, 224, 82, 430, 521]. varying-coefficient [589]. Vector [273, 306, 307, 235, 508, 438, 206]. vectors [412]. versus [291]. via [144, 576, 389, 428, 302, 483, 586, 588, 9, 11, 471, 305, 29]. view [355]. Visual [134]. Volatility [197, 527, 84]. Volume [37, 85].

W [409]. W-shaped [409]. Wald [398, 202]. warranty [274]. wavelet [50, 164, 551]. wavelet-based [551]. wavelets [479]. way [48, 257]. weak [492]. wealth [117]. webs [473]. Weibull [58, 150, 544, 248, 234, 375, 289, 339, 434, 143]. Weibull-model [234]. weight [418]. Weighted [321, 554, 422, 397, 586, 219, 232, 194, 343, 384, 382, 407]. weights [99, 219]. wet [205]. where [563]. which [233]. whose [607]. Win [396]. Win-probabilities [396]. Wishart [100]. withdrawal [317]. without [90]. working [43]. worse [402]. worst [402].

X [162]. X-ray [162]. XII [175].

Zelterman [512]. zero [570, 502, 115, 118]. zero-altered [115]. zero-inflated [570, 502]. zeros [442]. zip [332].

References

Balakrishnan:2004:E

[1] N. Balakrishnan. Editorial. Statistical Methodology, 1(1–2):v, December 2004. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                             //www.sciencedirect.com/science/
                                             article/pii/S1572312704000127.

Rao:2004:F

[2] C. R. Rao. Foreword. Statistical Methodology, 1(1–2):vii, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000115.

Casella:2004:MML

[3] George Casella, Christian P. Robert, and Martin T. Wells. Mixture models, latent variables and partitioned importance sampling. Statistical Methodology, 1(1–2):1–18, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000036.

Majumdar:2004:EDB

[4] Dibyen Majumdar and R. J. Martin. Efficient designs based on orthogonal arrays of type I and type II for experiments using units ordered over time or space. Statistical Methodology, 1 (1–2):19–35, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000097.

Street:2004:OSP

[5] Deborah J. Street and Leonie Burgess. Optimal stated preference choice experiments when all choice sets contain a specific option. Statistical Methodology, 1(1–2):37–45, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000048.

Ho:2004:PMP

[6] Clifford K. Ho. Probabilistic modeling of percutaneous absorption for risk-based exposure assessments and transdermal drug delivery. Statistical Methodology, 1(1–2):47–69, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270400005X.

Bagui:2004:CJU

[7] Subhash C. Bagui and Dulal K. Bhaumik. Computing Jacobians using exterior products. Statistical Methodology, 1(1–2):71–80, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000085.

Paindaveine:2004:UEP

[8] D. Paindaveine. A unified and elementary proof of serial and nonserial, univariate and multivariate, Chernoff– Savage results. Statistical Methodology, 1(1–2):81–91, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000073.

Lipovetsky:2004:TSB

[9] Stan Lipovetsky and W. Michael Conklin. Thurstone scaling via binary response regression. Statistical Methodology, 1(1–2):93–104, December 2004. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000024.

Angelidis:2004:UGM

[10] Timotheos Angelidis, Alexandros Benos, and Stavros Degiannakis. The use of GARCH models in VaR estimation. Statistical Methodology, 1(1– 2):105–128, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000103.

Salvadori:2004:BRP

[11] G. Salvadori. Bivariate return periods via 2-copulas. Statistical Methodology, 1(1–2):129–144, December 2004. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000061.

Anonymous:2004:AI

[12] Anonymous. Author index. Statistical Methodology, 1(1–2):145, December 2004. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S1572312704000206.

Cressie:2005:LBE

[13] Noel Cressie, Olivier Perrin, and Christine Thomas-Agnan. Likelihood-based estimation for Gaussian MRFs. Statistical Methodology, 2(1):1–16, March 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000139.

Lu:2005:DPN

[14] Yi Lu and José Garrido. Doubly periodic non-homogeneous Poisson models for hurricane data. Statistical Methodology, 2(1):17–35, March 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000140.

Lui:2005:SSC

[15] Kung-Jong Lui. Sample size calculation for testing non-inferiority and equivalence under Poisson distribution. Statistical Methodology, 2 (1):37–48, March 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312704000280.

Wilcox:2005:CSS

[16] Rand R. Wilcox. A comparison of six smoothers when there are multiple predictors. Statistical Methodology, 2(1):49–57, March 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312704000310.

Chan:2005:SAR

[17] Gary K. C. Chan and Hailiang Yang. Sensitivity analysis on ruin probabilities with heavy-tailed claims. Statistical Methodology, 2(1):59–63, March 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312704000279.

Datta:2005:EML

[18] Somnath Datta. Estimating the mean life time using right censored data. Statistical Methodology, 2 (1):65–69, March 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312704000309.

Delozier:2005:DIC

[19] Michael R. Delozier and Steven Orlich. Discovering influential cases in linear regression with MINITAB: Peeking into multidimensions with a MINITAB macro. Statistical Methodology, 2(2):71–81, July 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312704000292.

Magis:2005:IMD

[20] David Magis. On the influence of misclassified data on results of goodness of fit testing. Statistical Methodology, 2(2):82–94, July 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000043.

Huillet:2005:EPD

[21] Thierry Huillet and Christian Paroissin. Estimations of the parameter of a Dirichlet distribution using residual allocation model representations and sampling properties. Statistical Methodology, 2(2):95–110, July 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270400053X.

Taraldsen:2005:PEL

[22] Gunnar Taraldsen. A precise estimator for the log-normal mean. Statistical Methodology, 2(2):111–120, July 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270500002X.

Kumar:2005:EPS

[23] Somesh Kumar and Aditi Kar Gangopadhyay. Estimating parameters of a selected Pareto population. Statistical Methodology, 2(2):121–130, July 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000031.

Dudoit:2005:ACV

[24] Sandrine Dudoit and Mark J. van der Laan. Asymptotics of cross-validated risk estimation in estimator selection and performance assessment. Statistical Methodology, 2(2):131–154, July 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000158.

Lai:2005:CSP

[25] Dejian Lai. A comparative study on a permutation statistic. Statistical Methodology, 2(3):155–167, September 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S157231270500016X.

Cengiz:2005:BIB

[26] Mehmet A. Cengiz. Bayesian inference for bivariate generalized linear models in diagnosing renal arterial obstruction. Statistical Methodology, 2 (3):168–174, September 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000183.

Iwasaki:2005:BGL

[27] Masakazu Iwasaki and Hiroe Tsubaki. A bivariate generalized linear model with an application to meteorological data analysis. Statistical Methodology, 2(3):175–190, September 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000171.

Karunamuni:2005:BCK

[28] R. J. Karunamuni and T. Alberts. On boundary correction in kernel density estimation. Statistical Methodology, 2 (3):191–212, September 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000195.

Yen:2005:IRD

[29] Ester Yen and A. F. M. Smith. Image recognition via deformable templates. Statistical Methodology, 2 (3):213–225, September 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000213.

Hassairi:2005:BMC

[30] A. Hassairi and M. Zarai. Bhattacharyya matrices and cubic exponential families. Statistical Methodology, 2(3):226–232, September 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000225.

Gatto:2005:SAS

[31] Riccardo Gatto and Michael Mayer. Saddlepoint approximations for some models of circular data. Statistical Methodology, 2(4):233–248, December 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312705000201.

Davies:2005:CVM

[32] Simon L. Davies, Andrew A. Neath, and Joseph E. Cavanaugh. Cross validation model selection criteria for linear regression based on the Kullback– Leibler discrepancy. Statistical Methodology, 2(4):249–266, December 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000237.

Li:2005:ULL

[33] Chin-Shang Li. Using local linear kernel smoothers to test the lack of fit of nonlinear regression models. Statistical Methodology, 2(4): 267–284, December 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000341.

Fang:2005:IDM

[34] B. Q. Fang. Invariant distributions of the multivariate tests in the skew elliptical model. Statistical Methodology, 2(4):285–296, December 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000353.

Roy:2005:ITK

[35] Anuradha Roy and Ravindra Khattree. On implementation of a test for Kronecker product covariance structure for multivariate repeated measures data. Statistical Methodology, 2(4):297–306, December 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000389.

Bryant:2005:EMD

[36] Peter G. Bryant. An extended minimum description length (MDL) penalty function for binomial proportions. Statistical Methodology, 2 (4):307–315, December 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000481.

Anonymous:2005:AIV

[37] Anonymous. Author index volume 2 (2005). Statistical Methodology, 2 (4):317–318, December 2005. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000857.

Anonymous:2005:EB

[38] Anonymous. Editorial Board. Statistical Methodology, 2(4):CO2, December 2005. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000821.

Banks:2006:SIB

[39] David Banks and Grace S. Shieh. A special issue on bioinformatics: Introduction. Statistical Methodology, 3(1):vii–viii, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000717.

Anonymous:2006:CLS

[40] Anonymous. Contents list special issue. Statistical Methodology, 3(1):ix, January 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000924.

Imoto:2006:ETM

[41] Seiya Imoto, Tomoyuki Higuchi, Takao Goto, and Satoru Miyano. Error tolerant model for incorporating biological knowledge with expression data in estimating gene networks. Statistical Methodology, 3(1):1–16, January 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000675.

Aburatani:2006:GCM

[42] Sachiyo Aburatani, Shigeru Saito, Hiroyuki Toh, and Katsuhisa Horimoto. A graphical chain model for inferring regulatory system networks from gene expression profiles. Statistical Methodology, 3(1):17–28, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000778.

Zhu:2006:SBW

[43] X. Zhu, C. Ambroise, and G. J. McLachlan. Selection bias in working with the top genes in supervised classification of tissue samples. Statistical Methodology, 3(1):29–41, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000651.

Dyson:2006:MNP

[44] Greg Dyson and C. F. Jeff Wu. MAOSA: a new procedure for detection of differential gene expression. Statistical Methodology, 3(1):42–54, January 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270500047X.

Wang:2006:CDM

[45] Chuancai Wang, Jun Xie, and B. A. Craig. Context dependent models for discovery of transcription factor binding sites. Statistical Methodology, 3(1):55–68, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000663.

Boone:2006:AQT

[46] Edward L. Boone, Susan J. Simmons, Keying Ye, and Ann E. Stapleton. Analyzing quantitative trait loci for the Arabidopsis thaliana using Markov chain Monte Carlo model composition with restricted and unrestricted model spaces. Statistical Methodology, 3(1):69–78, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000687.

Datta:2006:FSM

[47] Susmita Datta and Lara M. De-Padilla. Feature selection and machine learning with mass spectrometry data for distinguishing cancer and noncancer samples. Statistical Methodology, 3(1):79–92, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270500064X.

Chandra:2006:NAI

[48] B. Chandra, S. Shanker, and Saroj Mishra. A new approach: Interrelated two-way clustering of gene expression data. Statistical Methodology, 3(1):93–102, January 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000699.

Anonymous:2006:EBa

[49] Anonymous. Editorial Board. Statistical Methodology, 3(1):CO2, January 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270500095X.

Mitra:2006:MER

[50] Sharmishtha Mitra and Amit Mitra. Modeling exchange rates using wavelet decomposed genetic neural networks. Statistical Methodology, 3(2):103–124, April 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000365.

Subramanian:2006:SAM

[51] Sundarraman Subramanian. Survival analysis for the missing censoring indicator model using kernel density estimation techniques. Statistical Methodology, 3(2):125–136, April 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000626.

Karr:2006:DQS

[52] Alan F. Karr, Ashish P. Sanil, and David L. Banks. Data quality: a statistical perspective. Statistical Methodology, 3(2):137–173, April 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000638.

Choudhury:2006:TPQ

[53] Gautam Choudhury and Madhuchanda Paul. A two phases queueing system with Bernoulli vacation schedule under multiple vacation policy. Statistical Methodology, 3(2):174–185, April 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000705.

Hayter:2006:SMC

[54] A. J. Hayter, H. P. Wynn, and W. Liu. Slope modified confidence bands for a simple linear regression model. Statistical Methodology, 3(2):186–192, April 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000729.

Anonymous:2006:EBb

[55] Anonymous. Editorial Board. Statistical Methodology, 3(2):CO2, April 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000062.

Penzer:2006:DSS

[56] Jeremy Penzer. Diagnosing seasonal shifts in time series using state space models. Statistical Methodology, 3(3):193–210, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000730.

Wilcox:2006:CRG

[57] Rand R. Wilcox. Comparing robust generalized variances and comments on efficiency. Statistical Methodology, 3(3):211–223, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000742.

Al-Saleh:2006:EWT

[58] Jamal A. Al-Saleh and Satish K. Agarwal. Extended Weibull type distribution and finite mixture of distributions. Statistical Methodology, 3 (3):224–233, July 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000754.

Bagui:2006:CMI

[59] Subhash Bagui, Sikha Bagui, A. Chatterjee, and K. L. Mehra. Classification with multiple independent measurements under a separate sampling scheme. Statistical Methodology, 3(3):234–251, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000766.

Tartakovsky:2006:DII

[60] Alexander G. Tartakovsky, Boris L. Rozovskii, Rudolf B. Blazek, and Hongjoong Kim. Detection of intrusions in information systems by sequential change-point methods. Statistical Methodology, 3(3):252–293, July 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000493. See

discussion [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71].

Gombay:2006:DDI

[61] Edit Gombay. A discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):294–295, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270500050X. See

[60].

Mukhopadhyay:2006:DDI

[62] Nitis Mukhopadhyay. A discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):296–298, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000511. See

[60].

Basseville:2006:DDI

[63] Michèle Basseville. A discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):299–303, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000523. See

[60].

Mei:2006:DDI

[64] Yajun Mei. A discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):304–306, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000535.

Zacks:2006:DDI

[65] Shelemyahu Zacks. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):307–309, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000547. See

[60].

Dragalin:2006:DDI

[66] Vladimir Dragalin. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):310–314, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000559. See

[60].

Kim:2006:DDI

[67] Seong-Hee Kim and James R. Wilson. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):315–319, July 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000560. See

[60].

Irle:2006:DDI

[68] Albrecht Irle. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):320–322, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000572. See

[60].

Chan:2006:DDI

[69] Hock Peng Chan. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):323–324, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000584. See

[60].

Pollak:2006:DDI

[70] Moshe Pollak. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):325–326, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000596. See

[60].

Khan:2006:DDI

[71] Rasul A. Khan. Discussion on “Detection of intrusions in information systems by sequential change-point methods” by Tartakovsky, Rozovskii, Blazek, and Kim. Statistical Methodology, 3(3):327–328, July 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000602. See

[60].

Tartakovsky:2006:AR

[72] Alexander G. Tartakovsky, Boris L. Rozovskii, Rudolf B. Blazek, and Hongjoong Kim. Authors’ response. Statistical Methodology, 3(3):329–340, July 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S1572312705000614.

Anonymous:2006:EBc

[73] Anonymous. Editorial Board. Statistical Methodology, 3(3):CO2, July 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000220.

Limnios:2006:NES

[74] Nikolaos Limnios and Brahim Ouhbi. Nonparametric estimation of some important indicators in reliability for semi-Markov processes. Statistical Methodology, 3(4):341–350, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270500078X.

Zimmerman:2006:TSE

[75] Donald W. Zimmerman. Two separate effects of variance heterogeneity on the validity and power of significance tests of location. Statistical Methodology, 3(4):351–374, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000791.

Ryabko:2006:UCB

[76] Boris Ryabko and Jaakko Astola. Universal codes as a basis for time series testing. Statistical Methodology, 3(4):375–397, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705000808.

Divanji:2006:LIL

[77] Gooty Divanji. Law of iterated logarithm and almost sure limit points for random sums. Statistical Methodology, 3(4):398–405, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705000936.

Lui:2006:SSC

[78] Kung-Jong Lui. Sample size calculation based on risk ratio under multiple matching. Statistical Methodology, 3(4):406–415, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705001036.

Rao:2006:MVI

[79] B. L. S. Prakasa Rao. Matrix variance inequalities for multivariate distributions. Statistical Methodology, 3(4):416–430, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312705001048.

Martin:2006:MEP

[80] R. J. Martin, T. Di Battista, L. Ippoliti, and E. Nissi. A model for estimating point sources in spatial data. Statistical Methodology, 3(4):431–443, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270500105X.

Lin:2006:QBL

[81] Lu Lin. Quasi Bayesian likelihood. Statistical Methodology, 3(4): 444–455, October 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705001061.

Walhin:2006:MPM

[82] J. F. Walhin and J. Paris. A mixed Poisson model with varying element sizes. Statistical Methodology, 3(4):456–463, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312705001073.

Soubeyrand:2006:RBS

[83] Samuel Soubeyrand, Joël Chadoeuf, Ivan Sache, and Christian Lannou. Residual-based specification of the random-effects distribution for cluster data. Statistical Methodology, 3(4):464–482, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000025.

Xu:2006:BEA

[84] Lixia Xu, Cihua Liu, and Gaoqin Nie. Bayesian estimation and the application of long memory stochastic volatility models. Statistical Methodology, 3(4):483–489, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000049.

Anonymous:2006:AIV

[85] Anonymous. Author index volume 3 (2006). Statistical Methodology, 3(4):491–494, October 2006. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000414.

Anonymous:2006:EBd

[86] Anonymous. Editorial Board. Statistical Methodology, 3(4):CO2, October 2006. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000384.

Bolfarine:2007:SPM

[87] Heleno Bolfarine and V. H. Lachos. Skew-probit measurement error models. Statistical Methodology, 4(1):1–12, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000037.

Pandit:2007:TEA

[88] Parameshwar V. Pandit and M. P. Anuradha. On testing exponentiality against new better than used of specified age. Statistical Methodology, 4(1):13–21, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000116.

Meintanis:2007:NGF

[89] Simos G. Meintanis. A new goodness-of-fit test for certain bivariate distributions applicable to traffic accidents. Statistical Methodology, 4(1):22–34, January 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000128.

Dryver:2007:ASR

[90] Arthur L. Dryver and Steven K. Thompson. Adaptive sampling without replacement of clusters. Statistical Methodology, 4(1):35–43, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270600013X.

Xu:2007:CAP

[91] Lei Xu, Sati Mazumdar, and Julie Price. Covariate adjustment in partial least squares (PLS) regression for the extraction of the spatial-temporal pattern from positron emission tomography data. Statistical Methodology, 4(1):44–63, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000323.

Neveu:2007:CAD

[92] D. Neveu, A. Kramar, and P. Dujols. Corrected asymptotic distribution of statistics based on the multinomial law. Statistical Methodology, 4(1):64–74, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000335.

Harel:2007:IMI

[93] Ofer Harel. Inferences on missing information under multiple imputation and two-stage multiple imputation. Statistical Methodology, 4 (1):75–89, January 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000153.

Kovtun:2007:LLS

[94] Mikhail Kovtun, Igor Akushevich, Kenneth G. Manton, and H. Dennis Tolley. Linear latent structure analysis: Mixture distribution models with linear constraints. Statistical Methodology, 4(1):90–110, January 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000177.

Mao:2007:EPS

[95] Chang Xuan Mao. Estimating population sizes by catch-effort methods. Statistical Methodology, 4(1): 111–119, January 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000189.

Anonymous:2007:EBa

[96] Anonymous. Editorial Board. Statistical Methodology, 4(1):CO2, January 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000566.

Subramanian:2007:MRA

[97] Sundarraman Subramanian. Median regression analysis from data with left and right censored observations. Statistical Methodology, 4(2):121–131, April 2007.

CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000141.

Magis:2007:DIS

[98] David Magis. On the detection of influential subsets of categories in goodness of fit testing. Statistical Methodology, 4(2):132–142, April 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000165.

Congdon:2007:MWM

[99] Peter Congdon. Model weights for model choice and averaging. Statistical Methodology, 4(2):143–157, April 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000190.

Sheena:2007:AEW

[100] Yo Sheena and Akimichi Takemura. An asymptotic expansion of Wishart distribution when the population eigenvalues are infinitely dispersed. Statistical Methodology, 4(2):158–184, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000207.

Kayid:2007:GFN

[101] M. Kayid. A general family of NBU classes of life distributions. Statistical Methodology, 4(2):185–195, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000347.

Fard:2007:FMA

[102] Mir Nabi Pirouzi Fard and Björn Holmquist. First moment approximations for order statistics from the extreme value distribution. Statistical Methodology, 4(2):196–203, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000359.

Hitchcock:2007:BBN

[103] David B. Hitchcock. Bandwidth-based nonparametric inference. Statistical Methodology, 4(2):204–216, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000360.

Gad:2007:SAL

[104] Ahmed M. Gad and Abeer S. Ahmed. Sensitivity analysis of longitudinal data with intermittent missing values. Statistical Methodology, 4 (2):217–226, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000463.

Choudhuri:2007:NBR

[105] Nidhan Choudhuri, Subhashis Ghosal, and Anindya Roy. Nonparametric binary regression using a Gaussian process prior. Statistical Methodology, 4(2):227–243, April 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312706000475.

Hassairi:2007:TDA

[106] A. Hassairi and O. Regaieg. On the Tukey depth of an atomic measure. Statistical Methodology, 4 (2):244–249, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000542.

Anonymous:2007:EBb

[107] Anonymous. Editorial board. Statistical Methodology, 4(2):IFC, April 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000111.

Beran:2007:CRM

[108] Jan Beran and Abdel-Yazid Karim Djaïdja. Credit risk modeling based on survival analysis with immunes. Statistical Methodology, 4(3):251–276, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S1572312706000487.

Kokonendji:2007:HDR

[109] Célestin C. Kokonendji, Clarice G. B.

Demétrio, and Silvio S. Zocchi. On Hinde–Demétrio regression models for overdispersed count data. Statistical Methodology, 4(3):277–291, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000499.

Lahiri:2007:NPR

[110] S. N. Lahiri, K. Furukawa, and Y.-D.

Lee. A nonparametric plug-in rule for selecting optimal block lengths for block bootstrap methods. Statistical Methodology, 4(3):292–321, July 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000505.

Zhang:2007:UED

[111] Biao Zhang. Using entropy distance measures for testing odds ratio parameters under logistic regression models based on case-control data. Statistical Methodology, 4(3):322–330, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000517.

Rousson:2007:SCB

[112] Valentin Rousson and Nicoleta Francisca Gosoniu. An R-square coefficient based on final prediction error. Statistical Methodology, 4(3):331–340, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000529.

Gatto:2007:GMD

[113] Riccardo Gatto and Sreenivasa Rao Jammalamadaka. The generalized von Mises distribution. Statistical Methodology, 4(3):341–353, July 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000530.

Liu:2007:HOP

[114] Junfeng Liu and Dipak K. Dey. Hierarchical overdispersed Poisson model with macrolevel autocorrelation. Statistical Methodology, 4(3):354–370, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000621.

Ghosh:2007:SIB

[115] Sujit K. Ghosh and Honggie Kim. Semiparametric inference based on a class of zero-altered distributions. Statistical Methodology, 4(3):371–383, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000020.

Anonymous:2007:EBc

[116] Anonymous. Editorial Board. Statistical Methodology, 4(3):IFC, July 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000251.

Kochar:2007:EWO

[117] Subhash Kochar, Xiaohu Li, and Maochao Xu. Excess wealth order and sample spacings. Statistical Methodology, 4(4):385–392, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000633.

Patil:2007:TPP

[118] M. K. Patil and D. T. Shirke. Testing parameter of the power series distribution of a zero inflated power series model. Statistical Methodology, 4(4):393–406, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312706000645.

Teerapabolarn:2007:BPB

[119] K. Teerapabolarn. A bound on the Poisson-binomial relative error. Statistical Methodology, 4(4):407–415, October 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S1572312706000657.

Mao:2007:TLE

[120] Chang Xuan Mao. Testing list effects and diagnosing individual catchabilities in the Rasch model. Statistical Methodology, 4(4):416–422, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000032.

Bandyopadhyay:2007:ANT

[121] Uttam Bandyopadhyay and Dhiman Dutta. Adaptive nonparametric tests for a single sample location problem. Statistical Methodology, 4(4): 423–433, October 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000044.

Bandyopadhyay:2007:TTA

[122] Uttam Bandyopadhyay and Saurav De. On a two-treatment adaptive allocation rule for continuous response. Statistical Methodology, 4 (4):434–447, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000056.

Rahman:2007:MSB

[123] M. Shafiqur Rahman and M. Ataharul Islam. Markov structure based logistic regression for repeated measures: an application to diabetes mellitus data. Statistical Methodology, 4(4):448–460, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000068.

Yonemoto:2007:ELE

[124] Koji Yonemoto and Takashi Yanagawa. Estimating the Lyapunov exponent from chaotic time series with dynamic noise. Statistical Methodology, 4(4):461–480, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000081.

Helmers:2007:NPE

[125] Roelof Helmers, I. Wayan Mangku, and Ricardas Zitikis. A non-parametric estimator for the doubly periodic Poisson intensity function. Statistical Methodology, 4(4):481–492, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000093.

Subramanian:2007:CMR

[126] Sundarraman Subramanian. Censored median regression and profile empirical likelihood. Statistical Methodology, 4(4):493–503, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000330.

Al-Saleh:2007:RLL

[127] Mohammad Fraiwan Al-Saleh. A rich learning lesson using the Poisson distribution. Statistical Methodology, 4(4):504–507, October 2007. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270700007X.

Schechtman:2007:TST

[128] Edna Schechtman and Michael Sherman. The two-sample t-test with a known ratio of variances. Statistical Methodology, 4(4):508–514, October 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312707000184.

Anonymous:2007:EBd

[129] Anonymous. Editorial Board. Statistical Methodology, 4(4):IFC, October 2007. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000457.

Beg:2008:CGO

[130] M. I. Beg and M. Ahsanullah. Concomitants of generalized order statistics from Farlie–Gumbel–Morgenstern distributions. Statistical Methodology, 5(1):1–20, January 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000214.

Choudhury:2008:NMG

[131] Gautam Choudhury. A note on the M^{x} /G/1 queue with a random setup time under a restricted admissibility policy with a Bernoulli vacation schedule. Statistical Methodology, 5(1):21–29, January 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000202.

Wang:2008:EPP

[132] Ji-Ping Wang and Bruce G. Lindsay. An exponential partial prior for improving nonparametric maximum likelihood estimation in mixture models. Statistical Methodology, 5(1):30–45, January 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000226.

Hutson:2008:ERR

[133] Alan D. Hutson. Estimating relative rank correlation: Theory and application. Statistical Methodology, 5 (1):46–55, January 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000238.

Weiss:2008:VAC

[134] Christian H. Weiß. Visual analysis of categorical time series. Statistical Methodology, 5(1):56–71, January 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000329.

Martins:2008:MOP

[135] A. P. Martins, H. Ferreira, and L. Pereira. Multidimensional outlierproneness of dependent data and the extremal index. Statistical Methodology, 5(1):72–82, January 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000378.

Malik:2008:CPD

[136] Malik Beshir Malik. Computational procedures for deriving sampling distributions of attribute-contaminated random variables. Statistical Methodology, 5(1):83–91, January 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000196.

Anonymous:2008:EBa

[137] Anonymous. Editorial Board. Statistical Methodology, 5(1):IFC, January 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270700069X.

Stavropoulos:2008:RTS

[138] A. Stavropoulos and C. Caroni. Rank test statistics for unbalanced nested designs. Statistical Methodology, 5 (2):93–105, March 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000342.

Sen:2008:CRE

[139] Arusharka Sen and Fang Tan. Curerate estimation under Case-1 interval censoring. Statistical Methodology, 5(2):106–118, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000354.

Hili:2008:HDE

[140] Ouagnina Hili. Hellinger distance estimation of general bilinear time series models. Statistical Methodology, 5(2):119–128, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000366.

Lui:2008:EPR

[141] Kung-Jong Lui. Estimation of proportion ratio in non-compliance randomized trials with repeated measurements in binary data. Statistical Methodology, 5(2):129–141, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270700038X.

Mathur:2008:ENT

[142] Sunil K. Mathur and Pamela F.

Smith. An efficient nonparametric test for bivariate two-sample location problem. Statistical Methodology, 5(2):142–159, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000391.

Wilding:2008:EAH

[143] Gregory E. Wilding and Govind S.

Mudholkar. Empirical approximations for Hoeffding’s test of bivariate independence using two Weibull extensions. Statistical Methodology, 5(2):160–170, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000408.

Ahsanullah:2008:CDR

[144] M. Ahsanullah and M. I. Beg. On characterizing distributions via regression of functions of generalized order statistics. Statistical Methodology, 5(2):171–176, March 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000512.

Hili:2008:EMT

[145] Ouagnina Hili. Estimation of a multiplethreshold AR(p) model. Statistical Methodology, 5(2):177–186, March 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000536.

Anonymous:2008:EBb

[146] Anonymous. Editorial Board. Statistical Methodology, 5(2):IFC, March 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000099.

Solvang:2008:CSA

[147] Hiroko Kato Solvang, Masanobu Taniguchi, Tomohiro Nakatani, and Shigeaki Amano. Classification and similarity analysis of fundamental frequency patterns in infant spoken language acquisition. Statistical Methodology, 5(3):187–208, May 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000500.

Buckley:2008:SCB

[148] Brooke E. Buckley and Walter W.

Piegorsch. Simultaneous confidence bands for Abbott-adjusted quantal response models in benchmark analysis. Statistical Methodology, 5 (3):209–219, May 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000524.

Grillenzoni:2008:DKS

[149] Carlo Grillenzoni. Design of kernel Msmoothers for spatial data. Statistical Methodology, 5(3):220–237, May 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000548.

Clarke:2008:BWT

[150] Jennifer Clarke and Mike West.

Bayesian Weibull tree models for survival analysis of clinico-genomic data. Statistical Methodology, 5 (3):238–262, May 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270700055X.

Nieto:2008:FUT

[151] Fabio H. Nieto. Forecasting with univariate TAR models. Statistical Methodology, 5(3):263–276, May 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000561.

Areia:2008:MSS

[152] Aníbal Areia, Manuela M. Oliveira, and João T. Mexia. Models for a series of studies based on geometrical representation. Statistical Methodology, 5(3):277–288, May 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312707000573.

Anonymous:2008:EBc

[153] Anonymous. Editorial Board. Statistical Methodology, 5(3):IFC, May 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000300.

Starck:2008:GE

[154] Jean-Luc Starck, Antoine Llebaria, and Tom Loredo. Guest editorial. Statistical Methodology, 5(4):v–vi, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000397.

Anonymous:2008:C

[155] Anonymous. CONTENTS. Statistical Methodology, 5(4):vii, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270800049X.

Abrial:2008:CDA

[156] P. Abrial, Y. Moudden, J.-L. Starck, J. Fadili, J. Delabrouille, and M. K. Nguyen. CMB data analysis and sparsity. Statistical Methodology, 5 (4):289–298, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000676.

Babu:2008:ODM

[157] G. Jogesh Babu, A. Mahabal, S. G.

Djorgovski, and R. Williams. Object detection in multi-epoch data. Statistical Methodology, 5(4):299–306, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000178.

Bobin:2008:SCR

[158] J. Bobin, Y. Moudden, J.-L. Starck, J. Fadili, and N. Aghanim. SZ and CMB reconstruction using generalized morphological component analysis. Statistical Methodology, 5 (4):307–317, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000627.

Bourguignon:2008:NMF

[159] Sébastien Bourguignon and Hervé Carfantan. New methods for fitting multiple sinusoids from irregularly sampled data. Statistical Methodology, 5(4):318–327, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000615.

Boursier:2008:DFM

[160] Y. Boursier and A. Llebaria. 3D dynamic forward modeling of polar plumes using hidden Markov trees. Statistical Methodology, 5(4):328–339, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270800004X.

Gissot:2008:OAU

[161] Samuel Gissot and J.-F. Hochedez. Oscillation analysis using optical flow: Application to an EUV movie of the solar corona. Statistical Methodology, 5(4):340–349, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000026.

Hojnacki:2008:UEC

[162] S. M. Hojnacki, G. Micela, S. M. LaLonde, E. D. Feigelson, and J. H. Kastner. An unsupervised, ensemble clustering algorithm: a new approach for classification of X-ray sources. Statistical Methodology, 5(4):350–360, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000233.

Jalobeanu:2008:MSD

[163] A. Jalobeanu, J. A. Gutiérrez, and E. Slezak. Multi-source data fusion and super-resolution from astronomical images. Statistical Methodology, 5(4):361–372, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000166.

Ordenovic:2008:DGS

[164] C. Ordénovic, C. Surace, B. Torrésani, and A. Llébaria. Detection of glitches and signal reconstruction using Hölder and wavelet analysis. Statistical Methodology, 5(4):373–386, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000038.

Zhang:2008:FPN

[165] B. Zhang, M. J. Fadili, J.-L. Starck, and S. W. Digel. Fast Poisson noise removal by biorthogonal Haar domain hypothesis testing. Statistical Methodology, 5(4):387–396, July 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000154.

Anonymous:2008:EBd

[166] Anonymous. Editorial Board. Statistical Methodology, 5(4):IFC, July 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000415.

Wilding:2008:SMT

[167] Gregory E. Wilding and Govind S.

Mudholkar. Some modifications of the Z-tests of normality and their isotones. Statistical Methodology, 5 (5):397–409, September 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000585.

Bohning:2008:SVF

[168] Dankmar Böhning. A simple variance formula for population size estimators by conditioning. Statistical Methodology, 5(5):410–423, September 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000597.

Mukhopadhyay:2008:TAN

[169] Nitis Mukhopadhyay and Basil M. de Silva. Theory and applications of a new methodology for the random sequential probability ratio test. Statistical Methodology, 5(5): 424–453, September 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000603.

Vellaisamy:2008:OCT

[170] P. Vellaisamy and M. Kumar. Optimal component test plans for a parallel system based on Type-II censoring. Statistical Methodology, 5 (5):454–461, September 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000639.

Karvanen:2008:EID

[171] Juha Karvanen. Efficient initial designs for binary response data. Statistical Methodology, 5(5):462–473, September 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312707000640.

Lui:2008:SSD

[172] Kung-Jong Lui and Chii-Dean Lin. Sample size determination for cluster randomization phase II trials of dichotomous data. Statistical Methodology, 5(5):474–485, September 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000664.

Anonymous:2008:EBe

[173] Anonymous. Editorial Board. Statistical Methodology, 5(5):IFC, September 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000555.

Siew:2008:GSL

[174] Hai-Yen Siew and Kunio Shimizu. The generalized symmetric Laplace distribution on the sphere. Statistical Methodology, 5(6):487–501, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000652.

Abd-Elfattah:2008:ESS

[175] A. M. Abd-Elfattah, Amal S. Hassan, and S. G. Nassr. Estimation in stepstress partially accelerated life tests for the Burr type XII distribution using type I censoring. Statistical Methodology, 5(6):502–514, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312707000755.

Subramanian:2008:HFE

[176] Sundarraman Subramanian and Derek Bean. Hazard function estimation from homogeneous right censored data with missing censoring indicators. Statistical Methodology, 5 (6):515–527, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000051.

Caroni:2008:TCP

[177] C. Caroni. Tests for censored paired data. Statistical Methodology, 5(6): 528–534, November 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000063.

Bandyopadhyay:2008:STP

[178] Uttam Bandyopadhyay, Amitava Mukherjee, and Barendra Purkait. Simultaneous tests for patterned recognition using nonparametric partially sequential procedure. Statistical Methodology, 5(6):535–551, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000075.

Congdon:2008:SAC

[179] Peter Congdon. A spatially adaptive conditional autoregressive prior for area health data. Statistical Methodology, 5(6):552–563, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000221.

Juutilainen:2008:MCV

[180] Ilmari Juutilainen and Juha Röning.

Modelling conditional variance function in industrial data: a case study. Statistical Methodology, 5 (6):564–575, November 2008. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000269.

Anonymous:2008:EBf

[181] Anonymous. Editorial Board. Statistical Methodology, 5(6):IFC, November 2008. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000671.

McElroy:2009:SDD

[182] Tucker McElroy and Scott Holan.

Spectral domain diagnostics for testing model proximity and disparity in time series data. Statistical Methodology, 6(1):1–20, January 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000142.

Boik:2009:PCC

[183] Robert J. Boik and Amin Shirvani. Principal components on coefficient of variation matrices. Statistical Methodology, 6(1):21–46, January 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000245.

Brua:2009:AEE

[184] J.-Y. Brua. Asymptotically efficient estimators for nonparametric heteroscedastic regression models. Statistical Methodology, 6(1):47–60, January 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000257.

Hwang:2009:LAN

[185] S. Y. Hwang, I. V. Basawa, and I. K. Yeo. Local asymptotic normality for bifurcating autoregressive processes and related asymptotic inference. Statistical Methodology, 6(1):61–69, January 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000270.

Jones:2009:KDB

[186] M. C. Jones. Kumaraswamy’s distribution: a beta-type distribution with some tractability advantages. Statistical Methodology, 6(1):70–81, January 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000282.

Nishikawa:2009:DAD

[187] Motoyoshi Nishikawa and Masanobu Taniguchi. Discriminant analysis for dynamics of stable processes. Statistical Methodology, 6(1):82–96, January 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270800035X.

Li:2009:AGP

[188] Jun Li, Daniel R. Jeske, and Jeff Pettyjohn. Approximate and generalized pivotal quantities for deriving confidence intervals for the offset between two clocks. Statistical Methodology, 6(1):97–107, January 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000361.

Anonymous:2009:EBa

[189] Anonymous. Editorial Board. Statistical Methodology, 6(1):IFC, January 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000750.

Adesola:2009:OED

[190] Oke Isaiah Adesola. Orthogonal experiments in the development of carbonresin for chloride removal from solutions. Statistical Methodology, 6 (2):109–119, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000373.

Dimitriou-Fakalou:2009:MDO

[191] Chrysoula Dimitriou-Fakalou. Modelling data observed irregularly over space and regularly in time. Statistical Methodology, 6(2):120–132, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000385.

Bhaumik:2009:HTP

[192] Dulal K. Bhaumik, Anindya Roy, Nicole A. Lazar, Kush Kapur, Subhash Aryal, John A. Sweeney, Dave Patterson, and Robert D. Gibbons. Hypothesis testing, power and sample size determination for between group comparisons in fMRI experiments. Statistical Methodology, 6(2):133–146, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000476.

Luong:2009:IPS

[193] Andrew Luong and Louis G. Doray. Inference for the positive stable laws based on a special quadratic distance. Statistical Methodology, 6 (2):147–156, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000488.

Vanbelle:2009:NLW

[194] S. Vanbelle and A. Albert. A note on the linearly weighted kappa coefficient for ordinal scales. Statistical Methodology, 6(2):157–163, March 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000506.

Lui:2009:ERR

[195] Kung-Jong Lui and Kuang-Chao Chang. Estimation of risk ratio in a noncompliance randomized clinical trial with trichotomous dose levels. Statistical Methodology, 6(2):164–176, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000518.

Krishna:2009:DBD

[196] Hare Krishna and Pramendra Singh Pundir. Discrete Burr and discrete Pareto distributions. Statistical Methodology, 6(2):177–188, March 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270800052X.

Kalaylioglu:2009:BUR

[197] Zeynep I. Kalaylioglu and Sujit K.

Ghosh. Bayesian unit-root tests for stochastic volatility models. Statistical Methodology, 6(2):189–201, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000531.

Beran:2009:ENR

[198] Jan Beran and Mark A. Heiler. Estimation of a nonparametric regression spectrum for multivariate time series. Statistical Methodology, 6 (2):202–222, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000610.

Anonymous:2009:EBb

[199] Anonymous. Editorial Board. Statistical Methodology, 6(2):IFC, March 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000069.

Mielke:2009:TRA

[200] P. W. Mielke, Jr., M. A. Long, K. J.

Berry, and J. E. Johnston. g-treatment ridit analyses: Resampling permutation methods. Statistical Methodology, 6(3):223–229, May 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000622.

Ngatchou-Wandji:2009:TSM

[201] Joseph Ngatchou-Wandji. Testing for symmetry in multivariate distributions. Statistical Methodology, 6 (3):230–250, May 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000634.

Li:2009:NIM

[202] Wenyun Li and Zehua Chen. A new interval mapping approach based on a general sib-pair regression model with a modified Wald test. Statistical Methodology, 6(3):251–261, May 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000646.

Alkasasbeh:2009:EGL

[203] Muna R. Alkasasbeh and Mohammad Z. Raqab. Estimation of the generalized logistic distribution parameters: Comparative study. Statistical Methodology, 6(3):262–279, May 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000658.

Jammalamadaka:2009:GCS

[204] S. Rao Jammalamadaka and Vasudevan Mangalam. A general censoring scheme for circular data. Statistical Methodology, 6(3):280–289, May 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000725.

Deni:2009:MPM

[205] Sayang Mohd Deni, Abdul Aziz Jemain, and Kamarulzaman Ibrahim. Mixed probability models for dry and wet spells. Statistical Methodology, 6(3):290–303, May 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000737.

Roy:2009:LBE

[206] Anindya Roy, Wayne A. Fuller, and YanYan Zhou. A likelihood based estimator for vector autoregressive processes. Statistical Methodology, 6 (3):304–319, May 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000816.

Kumon:2009:CEA

[207] Masayuki Kumon. On the conditions for the existence of ancillary statistics in a curved exponential family. Statistical Methodology, 6(3):320–335, May 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000828.

Anonymous:2009:EBc

[208] Anonymous. Editorial Board. Statistical Methodology, 6(3):IFC, May 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000112.

Cramer:2009:HIP

[209] Erhard Cramer. Hermite interpolation polynomials and distributions of ordered data. Statistical Methodology, 6(4):337–343, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270800083X.

Zografos:2009:FBG

[210] K. Zografos and N. Balakrishnan. On families of beta- and generalized gamma-generated distributions and associated inference. Statistical Methodology, 6(4):344–362, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000841.

Fauconnier:2009:ODM

[211] C. Fauconnier and G. Haesbroeck. Outliers detection with the minimum covariance determinant estimator in practice. Statistical Methodology, 6(4):363–379, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000853.

Koukouvinos:2009:GSM

[212] C. Koukouvinos and K. Mylona. Group screening method for the statistical analysis of E_{fNOD}-optimal mixed-level supersaturated designs. Statistical Methodology, 6(4):380–388, July 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312708000865.

Vellaisamy:2009:NUE

[213] P. Vellaisamy. A note on unbiased estimation following selection. Statistical Methodology, 6(4):389–396, July 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312708000877.

Samawi:2009:MPS

[214] Hani M. Samawi, Mohammad F. Al-Saleh, and Obaid Al-Saidy. The matched pair sign test using bivariate ranked set sampling for different ranking based schemes. Statistical Methodology, 6(4):397–407, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000033.

Patriota:2009:HSE

[215] A. G. Patriota, H. Bolfarine, and M. de Castro. A heteroscedastic structural errors-in-variables model with equation error. Statistical Methodology, 6(4):408–423, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000045.

Hobza:2009:RST

[216] Tomás Hobza, Domingo Morales, and Leandro Pardo. Rényi statistics for testing equality of autocorrelation coefficients. Statistical Methodology, 6(4):424–436, July 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000173.

Strauss:2009:R

[217] Linda Strauss and Durland Shumway. Review. Statistical Methodology, 6 (4):437–438, July 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000021.

Anonymous:2009:EBd

[218] Anonymous. Editorial Board. Statistical Methodology, 6(4):IFC, July 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000380.

Mielke:2009:NCW

[219] P. W. Mielke, Jr. and K. J. Berry. A note on Cohen’s weighted kappa coefficient of agreement with linear weights. Statistical Methodology, 6 (5):439–446, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000185.

Beran:2009:MDS

[220] Jan Beran, Yuanhua Feng, and Siegfried Heiler. Modifying the double smoothing bandwidth selector in nonparametric regression. Statistical Methodology, 6(5):447–465, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000197.

Soubeyrand:2009:ICB

[221] S. Soubeyrand, F. Carpentier, N. Desassis, and J. Chadoeuf. Inference with a contrast-based posterior distribution and application in spatial statistics. Statistical Methodology, 6 (5):466–477, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000203.

Islam:2009:ETL

[222] M. Ataharul Islam, Khalaf S. Sultan, and Rafiqul I. Chowdhury. Estimation and tests for a longitudinal regression model based on the Markov chain. Statistical Methodology, 6(5):478–489, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000227.

Felix-Medina:2009:LTS

[223] Martín H. Félix-Medina and Pedro E.

Monjardin. Link-tracing sampling with an initial sequential sample of sites: Estimating the size of a hidden human population. Statistical Methodology, 6(5):490–502, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000239.

Stefanescu:2009:LIE

[224] Catalina Stefanescu and Bruce W. Turnbull. Likelihood inference for exchangeable continuous data with covariates and varying cluster sizes; use of the Farlie–Gumbel–Morgenstern model. Statistical Methodology, 6 (5):503–512, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000240.

Timofeev:2009:NAS

[225] Andrey V. Timofeev. Non-asymptotic sequential confidence regions with fixed sizes for the multivariate nonlinear parameters of regression. Statistical Methodology, 6(5):513–526, September 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312709000252.

Lachos:2009:RMM

[226] V. H. Lachos, V. Garibay, F. V. Labra, and R. Aoki. A robust multivariate measurement error model with skew-normal/independent distributions and Bayesian MCMC implementation. Statistical Methodology, 6 (5):527–541, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000288.

Li:2009:UST

[227] Chin-Shang Li. Using P -splines to test the linearity of partially linear models. Statistical Methodology, 6 (5):542–552, September 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270900029X.

Anonymous:2009:EBe

[228] Anonymous. Editorial Board. Statistical Methodology, 6(5):IFC, September 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000495.

Fragiadakis:2009:TFN

[229] K. Fragiadakis, D. Karlis, and S. G.

Meintanis. Tests of fit for normal inverse Gaussian distributions. Statistical Methodology, 6(6):553–564, November 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312709000306.

Ando:2009:POS

[230] Tomohiro Ando. Penalized optimal scoring for the classification of multi-dimensional functional data. Statistical Methodology, 6(6):565–576, November 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312709000318.

Schneeweiss:2009:PRS

[231] H. Schneeweiss and J. Komlos. Probabilistic rounding and Sheppard’s correction. Statistical Methodology, 6 (6):577–593, November 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270900032X.

Subramanian:2009:ICW

[232] Sundarraman Subramanian and Gerhard Dikta. Inverse censoring weighted median regression. Statistical Methodology, 6(6):594–603, November 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000331.

Carter:2009:CTB

[233] Nancy J. Carter, Neil C. Schwertman, and Terry L. Kiser. A comparison of two boxplot methods for detecting univariate outliers which adjust for sample size and asymmetry. Statistical Methodology, 6(6):604–621, November 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000355.

Mudholkar:2009:TBF

[234] Govind S. Mudholkar, Kobby O. Asubonteng, and Alan D. Hutson. Transformation of the bathtub failure rate data in reliability for using Weibullmodel analysis. Statistical Methodology, 6(6):622–633, November 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000434.

Jamalizadeh:2009:EDL

[235] A. Jamalizadeh, H. Mahmoodian, and N. Balakrishnan. Exact distribution of a linear combination of a variable and order statistics from the other two variables of a trivariate elliptical random vector as a mixture of skew-elliptical distributions. Statistical Methodology, 6(6):634–644, November 2009. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000446.

Anonymous:2009:EBf

[236] Anonymous. Editorial Board. Statistical Methodology, 6(6):IFC, November 2009. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000720.

Bakouch:2010:HOM

[237] Hassan S. Bakouch. Higher-order moments, cumulants and spectral densities of the NGINAR(1) process. Statistical Methodology, 7(1):1–21, January 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000471.

Arashi:2010:ICI

[238] M. Arashi. Idea of constructing an image of Bayes action. Statistical Methodology, 7(1):22–29, January 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000458.

Nair:2010:CCD

[239] N. Unnikrishnan Nair and K. K. Sudheesh. Characterization of continuous distributions by properties of conditional variance. Statistical Methodology, 7(1):30–40, January 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270900046X.

Vila:2010:LOR

[240] Jean-Pierre Vila and Jean-Pierre Gauchi. Local optimality of replications from a minimal X-optimal design in regression: a sufficient and quasinecessary condition. Statistical Methodology, 7(1):41–57, January 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000562.

Al-Zahrani:2010:MIH

[241] Bander Al-Zahrani and Jordan Stoyanov. Moment inequalities for HN-BRUE with hypothesis testing application. Statistical Methodology, 7 (1):58–67, January 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000598.

Anonymous:2010:EBa

[242] Anonymous. Editorial Board. Statistical Methodology, 7(1):IFC, January 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000896.

Teerapabolarn:2010:PBR

[243] K. Teerapabolarn. A Poisson-binomial relative error uniform bound. Statistical Methodology, 7(2):69–76, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000604.

Afify:2010:EEP

[244] W. M. Afify. On estimation of the exponentiated Pareto distribution under different sample schemes. Statistical Methodology, 7(2):77–83, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000653.

Grafstrom:2010:EUP

[245] Anton Grafström. Entropy of unequal probability sampling designs. Statistical Methodology, 7(2):84–97, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000677.

Greer:2010:BIC

[246] Brandi A. Greer, Dean M. Young, and James D. Stamey. Bayesian inference for comparing two Poisson rates using data subject to underreporting and validation data. Statistical Methodology, 7(2):98–108, March 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000689.

Dharmasena:2010:FSR

[247] L. Sandamali Dharmasena and P. Zeephongsekul. Fitting software reliability growth curves using nonparametric regression methods. Statistical Methodology, 7(2):109–120, March 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000690.

Jazi:2010:DIW

[248] Mansour Aghababaei Jazi, Chin-Diew Lai, and Mohammad Hossein Alamatsaz. A discrete inverse Weibull distribution and estimation of its parameters. Statistical Methodology, 7 (2):121–132, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000707.

Al-Saleh:2010:EOU

[249] Mohammad Fraiwan Al-Saleh and Hani Samawi. On estimating the odds using moving extreme ranked set sampling. Statistical Methodology, 7(2):133–140, March 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000963.

Ndongo:2010:ELM

[250] Mor Ndongo, Abdou Kâ Diongue, Aliou Diop, and Simplice Dossou-Gbété. Estimation of long-memory parameters for seasonal fractional ARIMA with stable innovations. Statistical Methodology, 7(2):141–151, March 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000999.

Kumon:2010:SIQ

[251] Masayuki Kumon. Studies of information quantities and information geometry of higher order cumulant spaces. Statistical Methodology, 7 (2):152–172, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709001002.

Anonymous:2010:EBb

[252] Anonymous. Editorial Board. Statistical Methodology, 7(2):IFC, March 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000092.

Raftery:2010:SIS

[253] Adrian E. Raftery and Michael D.

Ward. Special issue on statistical methods for the social sciences: Guest editors’ introduction. Statistical Methodology, 7(3):173–174, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000067.

Fienberg:2010:PCS

[254] Stephen E. Fienberg. The prehistory of the center for statistics and the social sciences, with a prequel and epilogue. Statistical Methodology, 7(3):175–186, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000158.

Gelman:2010:BBD

[255] Andrew Gelman, Iwin Leenen, Iven Van Mechelen, Paul De Boeck, and Jeroen Poblome. Bridges between deterministic and probabilistic models for binary data. Statistical Methodology, 7(3):187–209, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000550.

Forster:2010:BIP

[256] Jonathan J. Forster. Bayesian inference for Poisson and multinomial loglinear models. Statistical Methodology, 7(3):210–224, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709001014.

Slavkovic:2010:STW

[257] Aleksandra B. Slavković and Juyoun Lee. Synthetic two-way contingency tables that preserve conditional frequencies. Statistical Methodology, 7(3):225–239, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231270900080X.

Dobra:2010:MOS

[258] Adrian Dobra and Héléne Massam. The mode oriented stochastic search (MOSS) algorithm for log-linear models with conjugate priors. Statistical Methodology, 7(3):240–253, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000215.

Jones:2010:DCR

[259] Bryan D. Jones, Chang-Jin Kim, and Richard Startz. Does Congress realign or smoothly adjust? A discrete switching model of Congressional partisan regimes. Statistical Methodology, 7(3):254–276, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000641.

Sriram:2010:CBM

[260] N. Sriram, Anthony G. Greenwald, and Brian A. Nosek. Correlational biases in mean response latency differences. Statistical Methodology, 7 (3):277–291, May 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000665.

Dobra:2010:MUM

[261] Adrian Dobra, Theo S. Eicher, and Alex Lenkoski. Modeling uncertainty in macroeconomic growth determinants using Gaussian graphical models. Statistical Methodology, 7(3):292–306, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000951.

Glynn:2010:EIS

[262] Adam N. Glynn and Jon Wakefield. Ecological inference in the social sciences. Statistical Methodology, 7(3):307–322, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000586.

Rudas:2010:IAC

[263] Tamás Rudas. Informative allocation and consistent treatment selection. Statistical Methodology, 7 (3):323–337, May 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000616.

Rubin:2010:DNM

[264] Donald B. Rubin and Elizabeth R. Zell.

Dealing with noncompliance and missing outcomes in a randomized trial using Bayesian technology: Prevention of perinatal sepsis clinical trial, Soweto, South Africa. Statistical Methodology, 7(3):338–350, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231270900063X.

Steele:2010:IMI

[265] Russell J. Steele, Naisyin Wang, and Adrian E. Raftery. Inference from multiple imputation for missing data using mixtures of normals. Statistical Methodology, 7(3):351–365, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000043.

Koskinen:2010:AER

[266] Johan H. Koskinen, Garry L. Robins, and Philippa E. Pattison. Analysing exponential random graph (p-star) models with missing data using Bayesian data augmentation. Statistical Methodology, 7(3):366–384, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000628.

Gormley:2010:MEL

[267] Isobel Claire Gormley and Thomas Brendan Murphy. A mixture of experts latent position cluster model for social network data. Statistical Methodology, 7(3):385–405, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000031.

Hays:2010:SMI

[268] Jude C. Hays, Aya Kachi, and Robert J.

Franzese, Jr. A spatial model incorporating dynamic, endogenous network interdependence: a political science application. Statistical Methodology, 7(3):406–428, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000975.

Sarkar:2010:PPU

[269] Avranil Sarkar, Stephen E. Fienberg, and David Krackhardt. Predicting profitability using advice branch bank networks. Statistical Methodology, 7(3):429–444, May 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000987.

Anonymous:2010:EBc

[270] Anonymous. Editorial Board. Statistical Methodology, 7(3):IFC, May 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000274.

Bobotas:2010:ENP

[271] Panayiotis Bobotas and Stavros Kourouklis. On the estimation of a normal precision and a normal variance ratio. Statistical Methodology, 7(4):445–463, July 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271000002X.

Kundu:2010:CAC

[272] Debasis Kundu and Rameshwar D. Gupta. A class of absolutely continuous bivariate distributions. Statistical Methodology, 7(4):464–477, July 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000055.

Patriota:2010:VAM

[273] Alexandre G. Patriota, João R. Sato, and Betsabé G. Blas Achic. Vector autoregressive models with measurement errors for testing Granger causality. Statistical Methodology, 7 (4):478–497, July 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271000016X.

Sharma:2010:EAS

[274] S. C. Sharma and Shinjita Agrawal. Effectiveness analysis of a system with a special warranty scheme. Statistical Methodology, 7(4):498–510, July 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000171.

Anonymous:2010:EBd

[275] Anonymous. Editorial Board. Statistical Methodology, 7(4):IFC, July 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000432.

Simpson:2010:ALR

[276] Sean L. Simpson. An adjusted likelihood ratio test for separability in unbalanced multivariate repeated measures data. Statistical Methodology, 7(5):511–519, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000183.

Islam:2010:PDS

[277] M. Ataharul Islam and Rafiqul I.

Chowdhury. Prediction of disease status: a regressive model approach for repeated measures. Statistical Methodology, 7(5):520–540, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000201.

Jensen:2010:LSE

[278] D. R. Jensen. Leverage and subset efficiencies in regression. Statistical Methodology, 7(5):541–551, September 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312710000407.

Hamid:2010:PLR

[279] Shahid Hamid, B. M. Golam Kibria, Sneh Gulati, Mark Powell, Bachir Annane, Steve Cocke, Jean-Paul Pinelli, Kurt Gurley, and Shu-Ching Chen. Predicting losses of residential structures in the state of Florida by the public hurricane loss evaluation model. Statistical Methodology, 7 (5):552–573, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000195. See

discussion [285, 280, 281, 282, 283, 284], response [286], and erratum [298].

Jagger:2010:DPH

[280] Thomas H. Jagger and James B. Elsner. Discussion on “Public Hurricane Loss Evaluation Models: Predicting losses of residential structures in the state of Florida” by S. Hamid et al. Statistical Methodology, 7 (5):574–576, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000663. See

[279].

Grigoriu:2010:DPH

[281] Mircea Grigoriu. Discussion on “Public hurricane loss evaluation models: Predicting losses of residential structures in the state of Florida” by S. Hamid et al. Statistical Methodology, 7(5):577–579, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000675. See

[279].

Johnson:2010:DPL

[282] Mark E. Johnson. Discussion of “Predicting losses of residential structures in the state of Florida by the Public Hurricane Loss Evaluation Models” by S. Hamid, B. M. G. Kibria, S. Gulati, M. Powell, B. Annane, S. Cocke, J.-P. Pinelli and S.- C. Chen. Statistical Methodology, 7 (5):580–582, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/

article/pii/S1572312710000687. See

[279].

El-Shaarawi:2010:DPH

[283] A. H. El-Shaarawi and Shojaeddin Chenouri. Discussion on “Public Hurricane Loss Evaluation Models: Predicting losses of residential structures in the state of Florida” by S. Hamid et al. Statistical Methodology, 7 (5):583–586, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000699. See

[279].

Pai:2010:DPH

[284] Jeffrey Pai. Discussion on “Public hurricane loss evaluation models: Predicting losses of residential structures in the state of Florida” by S. Hamid et al. Statistical Methodology, 7(5):587–591, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000705. See

[279].

Katz:2010:DPL

[285] Richard W. Katz. Discussion on “Predicting losses of residential structures in the state of Florida by the public hurricane loss evaluation model” by S. Hamid et al. Statistical Methodology, 7(5):592–595, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000717. See

Hamid:2010:ARD

[286] Shahid Hamid, B. M. Golam Kibria, Sneh Gulati, Mark Powell, Bachir Annane, Steve Cocke, Jean-Paul Pinelli, Kurt Gurley, and Shu-Ching Chen. Authors’ responses to the discussion on “Predicting losses of residential structures in the state of Florida by the Public Hurricane Loss Evaluation Model”. Statistical Methodology, 7 (5):596–600, September 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000729.

Anonymous:2010:EBe

[287] Anonymous. Editorial Board. Statistical Methodology, 7(5):IFC, September 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000572.

Tiwari:2010:TDO

[288] Neeraj Tiwari and A. K. Nigam. On two-dimensional optimal controlled nearest proportional to size sampling designs. Statistical Methodology, 7 (6):601–613, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000511.

Silva:2010:LWE

[289] Giovana Oliveira Silva, Edwin M. M. Ortega, and Vicente G. Cancho. Log-Weibull extended regression model: Estimation, sensitivity and residual analysis. Statistical Methodology, 7 (6):614–631, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000547.

Caroni:2010:TOP

[290] Chrys Caroni. Testing for outliers in the power-law growth parameters of sets of recurrent events data. Statistical Methodology, 7(6): 632–637, November 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000638.

Bar-Lev:2010:JVS

[291] Shaul K. Bar-Lev, Daoud Bshouty, Peter Grünwald, and Peter Harremoës. Jeffreys versus Shtarkov distributions associated with some natural exponential families. Statistical Methodology, 7(6):638–643, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271000064X.

Patriota:2010:IDM

[292] Alexandre G. Patriota, Artur J.

Lemonte, and Mário de Castro. Influence diagnostics in a multivariate normal regression model with general parameterization. Statistical Methodology, 7(6):644–654, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000651.

Wang:2010:MEP

[293] Dongliang Wang, Alan D. Hutson, and Jeffrey C. Miecznikowski. L-moment estimation for parametric survival models given censored data. Statistical Methodology, 7(6):655–667, November 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000808.

Rao:2010:PEB

[294] Poduri S. R. S. Rao. Performance of the empirical Bayes estimator for fixed parameters. Statistical Methodology, 7(6):668–672, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271000081X.

Warrens:2010:CKC

[295] Matthijs J. Warrens. Cohen’s kappa can always be increased and decreased by combining categories. Statistical Methodology, 7(6):673–677, November 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312710000535.

Iliopoulos:2010:OPS

[296] G. Iliopoulos and N. Balakrishnan. An odd property of sample median from odd sample sizes. Statistical Methodology, 7(6):678–686, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000821.

Dixit:2010:EEP

[297] U. J. Dixit and M. Jabbari Nooghabi. Efficient estimation in the Pareto distribution. Statistical Methodology, 7 (6):687–691, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000419.

Hamid:2010:EPL

[298] Shahid Hamid, B. M. Golam Kibria, Sneh Gulati, Mark Powell, Bachir Annane, Steve Cocke, Jean-Paul Pinelli, Kurt Gurley, and Shu-Ching Chen. Erratum to “Predicting losses of residential structures in the state of Florida by the public hurricane loss evaluation model [Stat. Methodol. 7 (2010) 552–573]. Statistical Methodology, 7(6):692, November 2010. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000791. See

[279].

Anonymous:2010:EBf

[299] Anonymous. Editorial Board. Statistical Methodology, 7(6):IFC, November 2010. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000882.

Banks:2011:ISI

[300] David Banks and Prem Goel. Introduction to a special issue on statistical data mining. Statistical Methodology, 8(1):1–2, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000833.

Asuncion:2011:ADE

[301] Arthur U. Asuncion, Padhraic Smyth, and Max Welling. Asynchronous distributed estimation of topic models for document analysis. Statistical Methodology, 8(1):3–17, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000213.

Buchman:2011:HDD

[302] Susan M. Buchman, Ann B. Lee, and Chad M. Schafer. High-dimensional density estimation via SCA: an example in the modelling of hurricane tracks. Statistical Methodology, 8(1):18–30, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000343.

Butucea:2011:QFE

[303] Cristina Butucea and Katia Meziani. Quadratic functional estimation in inverse problems. Statistical Methodology, 8(1):31–41, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000523.

Gunter:2011:VSQ

[304] L. Gunter, J. Zhu, and S. A. Murphy. Variable selection for qualitative interactions. Statistical Methodology, 8(1):42–55, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312709000276.

Wu:2011:NCL

[305] Yichao Wu and Yufeng Liu. Noncrossing large-margin probability estimation and its application to robust SVM via preconditioning. Statistical Methodology, 8(1):56–67, January 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000264.

Chakraborty:2011:BSS

[306] Sounak Chakraborty. Bayesian semisupervised learning with support vector machine. Statistical Methodology, 8(1):68–82, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312709000574.

Fokoue:2011:FBA

[307] Ernest Fokoué, Dongchu Sun, and Prem Goel. Fully Bayesian analysis of the relevance vector machine with an extended hierarchical prior structure. Statistical Methodology, 8(1):83–96, January 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000559.

Yu:2011:TPT

[308] Li Yu, Joseph S. Verducci, and Paul E.

Blower. The tau-path test for monotone association in an unspecified subpopulation: Application to chemogenomic data mining. Statistical Methodology, 8(1):97–111, January 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000079.

Anonymous:2011:EBa

[309] Anonymous. Editorial Board. Statistical Methodology, 8(1):IFC, January 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710001036.

Dhar:2011:SER

[310] Subhra Sankar Dhar and Probal Chaudhuri. On the statistical efficiency of robust estimators of multivariate location. Statistical Methodology, 8(2):113–128, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000845.

Gordon:2011:NOP

[311] Alexander Y. Gordon. A new optimality property of the Holm stepdown procedure. Statistical Methodology, 8(2):129–135, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000857.

Kakizawa:2011:NGB

[312] Yoshihide Kakizawa. A note on generalized Bernstein polynomial density estimators. Statistical Methodology, 8(2):136–153, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000869.

Ferreira:2011:SSM

[313] Clécio da Silva Ferreira, Heleno Bolfarine, and Víctor H. Lachos. Skew scale mixtures of normal distributions: Properties and estimation. Statistical Methodology, 8(2):154–171, March 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000948.

Economou:2011:MSC

[314] P. Economou. On model selection in the case of nested distributions — an application to frailty models. Statistical Methodology, 8(2):172–184, March 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271000095X.

Haridy:2011:UMA

[315] Salah Haridy, Zhang Wu, and Philippe Castagliola. Univariate and multivariate approaches for evaluating the capability of dynamic-behavior processes (case study). Statistical Methodology, 8(2):185–203, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000961.

Rahardja:2011:LBC

[316] Dewi Rahardja and Yan D. Zhao.

Likelihood-based confidence intervals for the risk difference of two-sample binary data with a fallible classifier and a gold standard. Statistical Methodology, 8(2):204–212, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000985.

Caroni:2011:TUW

[317] Chrys Caroni. Testing for undetected withdrawal in a follow-up study of times to events. Statistical Methodology, 8(2):213–220, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710000997.

Laib:2011:STC

[318] Naâmane Laïb and Fateh Chebana. A simultaneous test for conditional mean and conditional variance functions in time series models with martingale difference innovations. Statistical Methodology, 8(2):221–241, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710001000.

Ghoreishi:2011:DAM

[319] S. K. Ghoreishi and M. Alijani. Dynamic association modeling in 2×2 contingency tables. Statistical Methodology, 8(2):242–255, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710001012.

Gao:2011:INV

[320] Wei Gao, Zheng Tian, and Jiao R. Li. Identification of nonlinear VAR models using general conditional independence graphs. Statistical Methodology, 8(2):256–267, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312710001115.

Warrens:2011:WKH

[321] Matthijs J. Warrens. Weighted kappa is higher than Cohen’s kappa for tridiagonal agreement tables. Statistical Methodology, 8(2):268–272, March 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710000973.

Anonymous:2011:EBb

[322] Anonymous. Editorial Board. Statistical Methodology, 8(2):IFC, March 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710001164.

Zamba:2011:PEC

[323] K. D. Zamba and Akim Adekpedjou. Parameter estimation for correlated recurrent events under informative monitoring. Statistical Methodology, 8(3):273–290, May 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710001139.

Gillard:2011:AVC

[324] Jonathan Gillard. Asymptotic variancecovariance matrices for the linear structural model. Statistical Methodology, 8(3):291–303, May 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710001140.

Samawi:2011:THP

[325] Hani M. Samawi and Robert Vogel. Tests of homogeneity for partially matched-pairs data. Statistical Methodology, 8(3):304–313, May 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000037. See

corrigendum [393].

Strauss:2011:R

[326] Linda Strauss. Review. Statistical Methodology, 8(3):314–317, May 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000025.

Anonymous:2011:EBc

[327] Anonymous. Editorial Board. Statistical Methodology, 8(3):IFC, May 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000219.

Krivitsky:2011:ANS

[328] Pavel N. Krivitsky, Mark S. Handcock, and Martina Morris. Adjusting for network size and composition effects in exponential-family random graph models. Statistical Methodology, 8(4):319–339, July 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000086.

Dixit:2011:EEP

[329] U. J. Dixit and M. Jabbari Nooghabi. Efficient estimation in the Pareto distribution with the presence of outliers. Statistical Methodology, 8 (4):340–355, July 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000141.

Pepelyshev:2011:FDA

[330] Andrey Pepelyshev. Fixed-domain asymptotics of the maximum likelihood estimator and the Gaussian process approach for deterministic models. Statistical Methodology, 8 (4):356–362, July 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000177.

Chacko:2011:EPM

[331] Manoj Chacko and P. Yageen Thomas.

Estimation of parameter of Morgenstern type bivariate exponential distribution using concomitants of order statistics. Statistical Methodology, 8(4):363–376, July 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000189.

Kharrati-Kopaei:2011:IIP

[332] Mahmood Kharrati-Kopaei and Havva Faghih. Inferences for the inflation parameter in the zip distributions: the method of moments. Statistical Methodology, 8(4):377–388, July 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000190.

Lemonte:2011:MEM

[333] Artur J. Lemonte and Alexandre G.

Patriota. Multivariate elliptical models with general parameterization. Statistical Methodology, 8 (4):389–400, July 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100030X.

Petropoulos:2011:NCI

[334] Constantinos Petropoulos. New classes of improved confidence intervals for the scale parameter of a two-parameter exponential distribution. Statistical Methodology, 8(4):401–410, July 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000311.

Anonymous:2011:EBd

[335] Anonymous. Editorial Board. Statistical Methodology, 8(4):IFC, July 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000372.

dePascoa:2011:KGG

[336] Marcelino A. R. de Pascoa, Edwin M. M. Ortega, and Gauss M. Cordeiro. The Kumaraswamy generalized gamma distribution with application in survival analysis. Statistical Methodology, 8 (5):411–433, September 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000323.

Huang:2011:ADM

[337] Steward Huang, Ashis SenGupta, and Keh-Shin Lii. Application of Dirichlet mixture of normals in growth curve models. Statistical Methodology, 8 (5):434–441, September 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000359.

Lee:2011:DTA

[338] S. Lee and A. Karagrigoriou. A divergence test for autoregressive time series models. Statistical Methodology, 8(5):442–450, September 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100044X.

Soliman:2011:POW

[339] Ahmed A. Soliman, Abdullah Y. Al-Hossain, and Mashail M. Al-Harbi. Predicting observables from Weibull model based on general progressive censored data with asymmetric loss. Statistical Methodology, 8 (5):451–461, September 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000517.

Balakrishnan:2011:SCP

[340] N. Balakrishnan and J. P. Keating. On simultaneous closeness probabilities of order statistics from odd sample sizes to the population median. Statistical Methodology, 8(5): 462–467, September 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000505.

Thomas:2011:BR

[341] Hoben Thomas and James Martin. Book review. Statistical Methodology, 8(5):468–471, September 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000335.

Anonymous:2011:EBe

[342] Anonymous. Editorial Board. Statistical Methodology, 8(5):IFC, September 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000554.

Warrens:2011:CKW

[343] Matthijs J. Warrens. Cohen’s kappa is a weighted average. Statistical Methodology, 8(6):473–484, November 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000530.

Kundu:2011:EGE

[344] Debasis Kundu and Rameshwar D. Gupta. An extension of the generalized exponential distribution. Statistical Methodology, 8(6):485–496, November 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312711000657.

Kokonendji:2011:DAK

[345] Célestin C. Kokonendji and Tristan Senga Kiessé. Discrete associated kernels method and extensions. Statistical Methodology, 8 (6):497–516, November 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000670.

Weiss:2011:SCR

[346] Christian H. Weiß. Simultaneous confidence regions for the parameters of a Poisson INAR(1) model. Statistical Methodology, 8(6):517–527, November 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312711000682.

Peng:2011:NTS

[347] Liqian Peng and Tiejun Tong. A note on a two-sample T test with one variance unknown. Statistical Methodology, 8(6):528–534, November 2011. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000669.

Anonymous:2011:EBf

[348] Anonymous. Editorial Board. Statistical Methodology, 8(6):IFC, November 2011. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000797.

Murtagh:2012:ESI

[349] Fionn Murtagh, Tom Loredo, and Jean-Luc Starck. Editorial for the special issue on astrostatistics. Statistical Methodology, 9(1–2):1–3, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312711000724.

Dupe:2012:DUP

[350] F.-X. Dupé, M. J. Fadili, and J.-L. Starck. Deconvolution under Poisson noise using exact data fidelity and synthesis or analysis sparsity priors. Statistical Methodology, 9(1– 2):4–18, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000463.

Vega:2012:AIR

[351] Miguel Vega, Javier Mateos, Rafael Molina, and Aggelos K. Katsaggelos. Astronomical image restoration using variational methods and model combination. Statistical Methodology, 9(1– 2):19–31, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000347.

Bourguignon:2012:PMH

[352] Sébastien Bourguignon, David Mary, and Éric Slezak. Processing MUSE hyperspectral data: Denoising, deconvolution and detection of astrophysical sources. Statistical Methodology, 9(1– 2):32–43, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000487.

Petremand:2012:OBF

[353] M. Petremand, A. Jalobeanu, and C. Collet. Optimal Bayesian fusion of large hyperspectral astronomical observations. Statistical Methodology, 9(1– 2):44–54, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000451.

Bijaoui:2012:PEM

[354] A. Bijaoui, A. Recio-Blanco, P. de Laverny, and C. Ordénovic. Parameter estimation from a model grid application to the Gaia RVS spectra. Statistical Methodology, 9(1–2): 55–62, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000694.

Lachieze-Rey:2012:SSH

[355] Marc Lachièze-Rey. Spin spherical harmonics from Hopf fibration: a symplectic view. Statistical Methodology, 9(1– 2):63–70, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312710001127.

Pires:2012:TFM

[356] S. Pires, S. Plaszczynski, and A. Lavabre.

Towards a fast, model-independent cosmic microwave background bispectrum estimator. Statistical Methodology, 9(1– 2):71–84, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000748.

Labatie:2012:UPC

[357] Antoine Labatie, Jean-Luc Starck, Marc Lachièze-Rey, and Pablo Arnalte-Mur. Uncertainty in 2-point correlation function estimators and baryon acoustic oscillation detection in galaxy surveys. Statistical Methodology, 9(1–2): 85–100, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000499.

Loredo:2012:BMA

[358] Thomas J. Loredo, James O. Berger, David F. Chernoff, Merlise A. Clyde, and Bin Liu. Bayesian methods for analysis and adaptive scheduling of exoplanet observations. Statistical Methodology, 9(1–2):101–114, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312711000700.

Banerjee:2012:ESI

[359] Sudipto Banerjee and Dipak K. Dey. Editorial for the special issue on spatial statistics. Statistical Methodology, 9(1– 2):115–116, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000736.

Zhuang:2012:STM

[360] Lili Zhuang and Noel Cressie. Spatiotemporal modeling of sudden infant death syndrome data. Statistical Methodology, 9(1–2):117–143, January/March 2012. CODEN ????

ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312711000098.

Sahu:2012:CBM

[361] S. K. Sahu and K. S. Bakar.

A comparison of Bayesian models for daily ozone concentration levels. Statistical Methodology, 9(1–2): 144–157, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000475.

Debarsy:2012:IDS

[362] Nicolas Debarsy, Cem Ertur, and James P. LeSage. Interpreting dynamic space-time panel data models. Statistical Methodology, 9(1–2): 158–171, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000165.

Hoef:2012:PCE

[363] Jay M. Ver Hoef. Practical considerations for experimental designs of spatially autocorrelated data using computer intensive methods. Statistical Methodology, 9(1–2):172–184, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312711000438.

Ruiz-Cardenas:2012:EMC

[364] Ramiro Ruiz-Cárdenas, Marco A. R.

Ferreira, and Alexandra M. Schmidt. Evolutionary Markov chain Monte Carlo algorithms for optimal monitoring network designs. Statistical Methodology, 9(1–2):185–194, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000128.

Spock:2012:SSD

[365] Gunter Spöck and Ijaz Hussain. Spatial sampling design based on convex design ideas and using external drift variables for a rainfall monitoring network in Pakistan. Statistical Methodology, 9(1– 2):195–210, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000050.

Hitczenko:2012:STA

[366] M. Hitczenko and M. L. Stein. Some theory for anisotropic processes on the sphere. Statistical Methodology, 9(1– 2):211–227, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271100013X.

Song:2012:BMS

[367] Joon Jin Song and Victor De Oliveira. Bayesian model selection in spatial lattice models. Statistical Methodology, 9(1–2):228–238, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000049.

Zimmerman:2012:ESV

[368] Dale L. Zimmerman and Xiangming Fang. Estimating spatial variation in disease risk from locations coarsened by incomplete geocoding. Statistical Methodology, 9(1–2):239– 250, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000116.

Ghosh:2012:AUD

[369] Souparno Ghosh, Alan E. Gelfand, and Thomas Mølhave. Attaching uncertainty to deterministic spatial interpolations. Statistical Methodology, 9(1– 2):251–264, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000529.

Reich:2012:NBM

[370] Brian J. Reich and Montserrat Fuentes.

Nonparametric Bayesian models for a spatial covariance. Statistical Methodology, 9(1–2):265–274, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312711000104.

Lemos:2012:CLM

[371] Ricardo T. Lemos and Bruno Sansó.

Conditionally linear models for nonhomogeneous spatial random fields. Statistical Methodology, 9(1–2):275– 284, January/March 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000153.

Anonymous:2012:EBa

[372] Anonymous. Editorial Board. Statistical Methodology, 9(1–2):IFC, January/March 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312711000980.

Mukhopadhyay:2012:CDG

[373] S. Mukhopadhyay and A. I. Khuri. Comparison of designs for generalized linear models under model misspecification. Statistical Methodology, 9(3):285–304, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271100075X.

Basak:2012:ETP

[374] Indrani Basak and N. Balakrishnan. Estimation for the three-parameter gamma distribution based on progressively censored data. Statistical Methodology, 9(3):305–319, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000761.

Qian:2012:FIM

[375] Lianfen Qian. The Fisher information matrix for a three-parameter exponentiated Weibull distribution under type II censoring. Statistical Methodology, 9(3):320–329, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000876.

Warrens:2012:FMR

[376] Matthijs J. Warrens. A family of multi-rater kappas that can always be increased and decreased by combining categories. Statistical Methodology, 9(3):330–340, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000888.

Warrens:2012:ECC

[377] Matthijs J. Warrens. The effect of combining categories on Bennett, Alpert and Goldstein’s S. Statistical Methodology, 9(3):341–352, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100089X.

Nkurunziza:2012:CES

[378] Sévérien Nkurunziza. Constrained estimation and some useful results in several multivariate models. Statistical Methodology, 9(3):353–363, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000906.

Jayasinghe:2012:NSE

[379] Chathuri L. Jayasinghe and P. Zeephongsekul. On the nonparametric smooth estimation of the reversed hazard rate function. Statistical Methodology, 9(3):364–380, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000918.

Rastogi:2012:EPB

[380] Manoj Kumar Rastogi and Yogesh Mani Tripathi. Estimating the parameters of a Burr distribution under progressive type II censoring. Statistical Methodology, 9(3):381–391, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000931.

Takemura:2012:MCM

[381] Akimichi Takemura and Hisayuki Hara. Markov chain Monte Carlo test of toric homogeneous Markov chains. Statistical Methodology, 9(3):392–406, May 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711000955.

Warrens:2012:EWK

[382] Matthijs J. Warrens. Equivalences of weighted kappas for multiple raters. Statistical Methodology, 9 (3):407–422, May 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711001171.

Withers:2012:TME

[383] Christopher S. Withers and Saralees Nadarajah. Transformations of multivariate Edgeworth type expansions. Statistical Methodology, 9 (3):423–439, May 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312711001183.

Warrens:2012:CQW

[384] Matthijs J. Warrens. Cohen’s quadratically weighted kappa is higher than linearly weighted kappa for tridiagonal agreement tables. Statistical Methodology, 9(3):440–444, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000773.

Hajebi:2012:CIG

[385] Mahtab Hajebi, Sadegh Rezaei, and Saralees Nadarajah. Confidence intervals for P (Y < X) for the generalized exponential distribution. Statistical Methodology, 9(3):445–455, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100092X.

Davidov:2012:OIR

[386] Ori Davidov. Ordered inference, rank statistics and combining p-values: a new perspective. Statistical Methodology, 9(3):456–465, May 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711000943.

Anonymous:2012:EBb

[387] Anonymous. Editorial Board. Statistical Methodology, 9(3):IFC, May 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711001225.

Brahimi:2012:SEC

[388] Brahim Brahimi and Abdelhakim Necir.

A semiparametric estimation of copula models based on the method of moments. Statistical Methodology, 9(4):467–477, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711001195.

Balakrishnan:2012:ASD

[389] N. Balakrishnan, C. Koukouvinos, and C. Parpoula. Analysis of a supersaturated design using entropy prior complexity for binary responses via generalized linear models. Statistical Methodology, 9(4):478–485, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100116X.

Malik:2012:EDS

[390] Malik Beshir Malik. Extensions and developments on the Schilling and Dodge mixed dependent acceptance sampling plans. Statistical Methodology, 9(4):486–489, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711001201.

Lange:2012:SSR

[391] Katharina Lange and Edgar Brunner. Sensitivity, specificity and ROCcurves in multiple reader diagnostic trials — a unified, nonparametric approach. Statistical Methodology, 9(4):490–500, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312711001341.

Sankaran:2012:NES

[392] P. G. Sankaran, M. Manoharan, and P. Anisha. Nonparametric estimators for a survivor function of paired recurrent events. Statistical Methodology, 9(4):501–512, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000020.

Samawi:2012:CTH

[393] Hani M. Samawi, Robert Vogel, and Wilson Koech. Corrigendum to “Tests of homogeneity for partially matched-pairs data” [Stat. Methodol. 8 (2011) 304–313]. Statistical Methodology, 9(4):513, July 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271100133X. See

[325].

Anonymous:2012:EBc

[394] Anonymous. Editorial Board. Statistical Methodology, 9(4):IFC, July 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000068.

Rao:2012:EBL

[395] Poduri S. R. S. Rao. Estimation of the best linear unbiased predictor for the mean with unequal sample sizes. Statistical Methodology, 9 (5):515–519, September 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000032.

Hayter:2012:WPR

[396] A. J. Hayter. Win-probabilities for regression models. Statistical Methodology, 9(5):520–527, September 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000044.

Hu:2012:SLL

[397] Xueping Hu and Guohua Fang. Strong laws of large numbers and mean convergence theorems for randomly weighted sums of arrays under a condition of integrability. Statistical Methodology, 9(5):528–536, September 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000147.

Lemonte:2012:LPS

[398] Artur J. Lemonte and Silvia L. P. Ferrari. Local power and size properties of the LR, Wald, score and gradient tests in dispersion models. Statistical Methodology, 9(5):537–554, September 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000160.

Spanos:2012:RBL

[399] Aris Spanos. Revisiting the Berger location model: Fallacious confidence interval or a rigged example? Statistical Methodology, 9 (5):555–561, September 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000159.

Anonymous:2012:EBd

[400] Anonymous. Editorial Board. Statistical Methodology, 9(5):IFC, September 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000226.

Volterman:2012:PCC

[401] William Volterman, Katherine F.

Davies, and N. Balakrishnan. Pitman closeness as a criterion for the determination of the optimal progressive censoring scheme. Statistical Methodology, 9(6):563–572, November 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000196.

Gill:2012:SST

[402] Amar Nath Gill, Anju Goyal, and Parminder Singh. Simultaneous selection of treatments better and worse than the best and worst controls. Statistical Methodology, 9(6): 573–588, November 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000172.

Alzaatreh:2012:DAC

[403] Ayman Alzaatreh, Carl Lee, and Felix Famoye. On the discrete analogues of continuous distributions. Statistical Methodology, 9 (6):589–603, November 2012. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000184.

Wang:2012:AMT

[404] Min zhen Wang and Kunio Shimizu. On applying Möbius transformation to cardioid random variables. Statistical Methodology, 9(6):604–614, November 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312712000202.

Anonymous:2012:EBe

[405] Anonymous. Editorial Board. Statistical Methodology, 9(6):IFC, November 2012. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000329.

Adekpedjou:2013:SIC

[406] Akim Adekpedjou, Jonathan Quiton, and Xuerong Meggie Wen. Semiparametric inference with correlated recurrence time data. Statistical Methodology, 10(1):1–13, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000287.

Warrens:2013:CIB

[407] Matthijs J. Warrens. Conditional inequalities between Cohen’s kappa and weighted kappas. Statistical Methodology, 10(1):14–22, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271200038X.

Marra:2013:VSB

[408] Giampiero Marra. On p-values for semiparametric bivariate probit models. Statistical Methodology, 10(1): 23–28, January 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000305.

Gallop:2013:MES

[409] Robert J. Gallop, Randall H. Rieger, Scott McClintock, and David C. Atkins. A model for extreme stacking of data at endpoints of a distribution: Illustration with W-shaped data. Statistical Methodology, 10(1):29–45, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000299.

Bean:2013:FSM

[410] Gordon J. Bean, Elizabeth A. Dimarco, Laina D. Mercer, Laura K. Thayer, Anindya Roy, and Subhashis Ghosal. Finite skew-mixture models for estimation of positive false discovery rates. Statistical Methodology, 10(1):46–57, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000391.

Rezapour:2013:PPT

[411] M. Rezapour, M. H. Alamatsaz, N. Balakrishnan, and E. Cramer. On properties of progressively Type-II censored order statistics arising from dependent and non-identical random variables. Statistical Methodology, 10 (1):58–71, January 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000408.

Mathur:2013:NSR

[412] Sunil Mathur and Mohammad B. Sepehrifar. A new signed rank test based on slopes of vectors for bivariate location problems. Statistical Methodology, 10(1):72–84, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271200041X.

Gaur:2013:NTM

[413] Anil Gaur, Kalpana K. Mahajan, and Sangeeta Arora. A nonparametric test for a multi-sample scale problem using ranked-set data. Statistical Methodology, 10(1):85–92, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000421.

Nair:2013:SNA

[414] N. Unnikrishnan Nair and P. G. Sankaran. Some new applications of the total time on test transforms. Statistical Methodology, 10 (1):93–102, January 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000433.

Bose:2013:NCR

[415] Arup Bose and Sreela Gangopadhyay. A note on concomitants of records. Statistical Methodology, 10 (1):103–112, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000445.

Wang:2013:DDA

[416] Fangpo Wang and Alan E. Gelfand.

Directional data analysis under the general projected normal distribution. Statistical Methodology, 10 (1):113–127, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000457.

Cramer:2013:SED

[417] Erhard Cramer and N. Balakrishnan.

On some exact distributional results based on Type-I progressively hybrid censored data from exponential distributions. Statistical Methodology, 10(1):128–150, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000469.

Autin:2013:TMS

[418] F. Autin and C. Pouet. Testing the means of subgroups in the varying mixing weight model. Statistical Methodology, 10(1):151–169, January 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000470.

Anonymous:2013:EBa

[419] Anonymous. Editorial Board. Statistical Methodology, 10(1):IFC, January 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000512.

Nandram:2013:BPI

[420] Balgobin Nandram, Dilli Bhatta, Dhiman Bhadra, and Gang Shen. Bayesian predictive inference of a finite population proportion under selection bias. Statistical Methodology, 11(??):1–21, March 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000494.

Bhandary:2013:ATE

[421] Madhusudan Bhandary and Hongying Dai. An alternative test for the equality of variances for several populations in randomised complete block design. Statistical Methodology, 11(??):22–35, March 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000482.

Alvarez-Andrade:2013:SAW

[422] Sergio Alvarez-Andrade and Salim Bouzebda. Strong approximations for weighted bootstrap of empirical and quantile processes with applications. Statistical Methodology, 11 (??):36–52, March 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000615.

Han:2013:CAI

[423] Bing Han. Conditional Akaike information criterion in the Fay–Herriot model. Statistical Methodology, 11 (??):53–67, March 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000627.

Sirdari:2013:STM

[424] Mahboobeh Zangeneh Sirdari, M. Ataharul Islam, and Norhashidah Awang. A stationarity test on Markov chain models based on marginal distribution. Statistical Methodology, 11 (??):68–76, March 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000639.

Anonymous:2013:EBb

[425] Anonymous. Editorial Board. Statistical Methodology, 11(??):IFC, March 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000688.

Withers:2013:CFE

[426] Christopher S. Withers and Saralees Nadarajah. Cornish–Fisher expansions for functionals of the partial sum empirical distribution. Statistical Methodology, 12(??):1–15, May 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000767.

Withers:2013:EDA

[427] Christopher S. Withers and Saralees Nadarajah. Expansions for the distribution of asymptotically chi-square statistics. Statistical Methodology, 12(??):16–30, May 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000652.

Beg:2013:CRG

[428] M. I. Beg, M. Ahsanullah, and Ramesh C. Gupta. Characterizations via regression of generalized order statistics. Statistical Methodology, 12(??):31–41, May 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000779.

Bandyopadhyay:2013:TCO

[429] Uttam Bandyopadhyay and Shirsendu Mukherjee. Test for conditional odds ratio in matching pairs inverse sampling design. Statistical Methodology, 12(??):42–59, May 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000664.

Zhao:2013:MSV

[430] Peixin Zhao and Gaorong Li. Modified SEE variable selection for varying coefficient instrumental variable models. Statistical Methodology, 12 (??):60–70, May 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000780.

Bhattacharya:2013:FBA

[431] Sourabh Bhattacharya. A fully Bayesian approach to assessment of model adequacy in inverse problems. Statistical Methodology, 12(??):71–83, May 2013.

CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000755.

Anonymous:2013:EBc

[432] Anonymous. Editorial Board. Statistical Methodology, 12(??):IFC, May 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271300004X.

Scollnik:2013:CTP

[433] David P. M. Scollnik. Comments on two papers concerning estimation of the parameters of the Pareto distribution in the presence of outliers. Statistical Methodology, 13(??):1–11, July 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000822.

Teimouri:2013:BCM

[434] Mahdi Teimouri and Saralees Nadarajah. Bias corrected MLEs for the Weibull distribution based on records. Statistical Methodology, 13(??):12–24, July 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000026.

Roy:2013:PPB

[435] Vivekananda Roy and Mark S. Kaiser.

Posterior propriety for Bayesian binomial regression models with a parametric family of link functions. Statistical Methodology, 13(??):25–41, July 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312712000834.

Harris:2013:SAL

[436] Ian R. Harris. A simple approximation to the likelihood interval for a binomial proportion. Statistical Methodology, 13(??):42–47, July 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000130.

Cancho:2013:DNB

[437] Vicente G. Cancho, Dipankar Bandyopadhyay, Francisco Louzada, and Bao Yiqi. The destructive negative binomial cure rate model with a latent activation scheme. Statistical Methodology, 13(??):48–68, July 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000142.

Nagakura:2013:EVE

[438] Daisuke Nagakura. Explicit vector expression of exact score for time series models in state space form. Statistical Methodology, 13(??):69–74, July 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000117.

Anonymous:2013:EBd

[439] Anonymous. Editorial Board. Statistical Methodology, 13(??):IFC, July 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000191.

Afroughi:2013:MAS

[440] Solaiman Afroughi, Mehdi Ghandehari Motlagh, Soghrat Faghihzadeh, and Majid Jafari Khaledi. A model for analyzing spatially correlated binary data clustered in uncorrelated lattices. Statistical Methodology, 14 (??):1–14, September 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000129.

Qin:2013:MSG

[441] Xu Qin, Jiang-She Zhang, and Xiao-Dong Yan. Model selection of the generalized von Mises distribution based on empirical mode decomposition with data analyses. Statistical Methodology, 14(??):15–25, September 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000154.

Yang:2013:MRM

[442] Ming Yang, Gideon K. D. Zamba, and Joseph E. Cavanaugh. Markov regression models for count time series with excess zeros: a partial likelihood approach. Statistical Methodology, 14(??):26–38, September 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000166.

Bouzebda:2013:SML

[443] Salim Bouzebda and Amor Keziou. A semiparametric maximum likelihood ratio test for the change point in copula models. Statistical Methodology, 14(??):39–61, September 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000270.

Mukhopadhyay:2013:RBI

[444] Nitis Mukhopadhyay and Mun S. Son. Ratios X/Z, Y/Z built from independent random variables (X, Y) and Z may not always be dependent. Statistical Methodology, 14 (??):62–66, September 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000178.

Anonymous:2013:EBe

[445] Anonymous. Editorial Board. Statistical Methodology, 14(??):IFC, September 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312713000324.

Midhu:2013:CDL

[446] N. N. Midhu, P. G. Sankaran, and N. Unnikrishnan Nair. A class of distributions with the linear mean residual quantile function and its generalizations. Statistical Methodology, 15(??):1–24, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000294.

Berrocal:2013:SCD

[447] Veronica J. Berrocal, Marie Lynn Miranda, Alan E. Gelfand, and Sourabh Bhattacharya. Synthesizing categorical datasets to enhance inference. Statistical Methodology, 15 (??):25–45, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000300.

Lv:2013:SEL

[448] Xiaofeng Lv and Rui Li. Smoothed empirical likelihood confidence intervals for quantile regression parameters with auxiliary information. Statistical Methodology, 15(??):46–54, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000403.

Rezapour:2013:EBT

[449] M. Rezapour and N. Balakrishnan. Estimators based on trimmed Kendall’s tau in multivariate copula models. Statistical Methodology, 15 (??):55–72, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000452.

Asgharzadeh:2013:ESS

[450] A. Asgharzadeh, R. Valiollahi, and Mohammad Z. Raqab. Estimation of the stress-strength reliability for the generalized logistic distribution. Statistical Methodology, 15 (??):73–94, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000464.

Mitra:2013:SSS

[451] S. Mitra, A. Ganguly, D. Samanta, and D. Kundu. On the simple step-stress model for two-parameter exponential distribution. Statistical Methodology, 15(??):95–114, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000415.

Ryabko:2013:CSA

[452] Boris Ryabko and Daniil Ryabko. A confidence-set approach to signal denoising. Statistical Methodology, 15 (??):115–120, November 2013. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000476.

Anonymous:2013:EBf

[453] Anonymous. Editorial Board. Statistical Methodology, 15(??):IFC, November 2013. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312713000506.

Ghosh:2014:PCC

[454] Indranil Ghosh and N. Balakrishnan. Partial or complete characterization of a bivariate distribution based on one conditional distribution and partial specification of the mode function of the other conditional distribution. Statistical Methodology, 16(??):1–13, January 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000567.

Ibacache-Pulgar:2014:IDE

[455] Germán Ibacache-Pulgar, Gilberto A.

Paula, and Manuel Galea. On influence diagnostics in elliptical multivariate regression models with equicorrelated random errors. Statistical Methodology, 16(??):14–31, January 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000488.

Crawford:2014:EER

[456] Scott D. Crawford. Efficient estimation in a regression model with missing responses. Statistical Methodology, 16(??):32–46, January 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000579.

Barbiero:2014:ADS

[457] Alessandro Barbiero. An alternative discrete skew Laplace distribution. Statistical Methodology, 16 (??):47–67, January 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000580.

Chan:2014:NAI

[458] Kwun Chuen Gary Chan. A note about the identifiability of causal effect estimates in randomized trials with non-compliance. Statistical Methodology, 16(??):68–71, January 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000592.

Rajesh:2014:BED

[459] G. Rajesh, E. I. Abdul-Sathar, K. R. Muraleedharan Nair, and K. V. Reshmi. Bivariate extension of dynamic cumulative residual entropy. Statistical Methodology, 16(??):72–82, January 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000646.

Balakrishnan:2014:UMP

[460] N. Balakrishnan, Ghobad Barmalzan, and Abedin Haidari. Uniformly most powerful test with two-dimensional minimal sufficient statistic. Statistical Methodology, 16(??):83–89, January 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312713000610.

Dutta:2014:MCM

[461] Somak Dutta and Sourabh Bhattacharya. Markov chain Monte Carlo based on deterministic transformations. Statistical Methodology, 16 (??):100–116, January 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000683.

Anonymous:2014:EBa

[462] Anonymous. Editorial Board. Statistical Methodology, 16(??):IFC, January 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000762.

Anonymous:2014:E

[463] Anonymous. Editorial. Statistical Methodology, 17(??):1–2, March 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000932.

Xu:2014:GHJ

[464] Yuan Xu, Laura L. E. Cowen, and Caleb Gardner. Group heterogeneity in the Jolly–Seber-tag-loss model. Statistical Methodology, 17(??):3–16, March 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000105.

Hubbard:2014:PRC

[465] B. A. Hubbard, D. J. Cole, and B. J. T. Morgan. Parameter redundancy in capture-recapture-recovery models. Statistical Methodology, 17 (??):17–29, March 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000809.

King:2014:GLF

[466] R. King and R. S. McCrea. A generalised likelihood framework for partially observed capture-recapture-recovery models. Statistical Methodology, 17(??):30–45, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000609.

Hoef:2014:SHM

[467] Jay M. Ver Hoef, Michael F. Cameron, Peter L. Boveng, Josh M. London, and Erin E. Moreland. A spatial hierarchical model for abundance of three ice-associated seal species in the eastern Bering Sea. Statistical Methodology, 17(??):46–66, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000282.

Zipkin:2014:FSD

[468] Elise F. Zipkin, Jeffery B. Leirness, Brian P. Kinlan, Allan F. O’Connell, and Emily D. Silverman. Fitting statistical distributions to sea duck count data: Implications for survey design and abundance estimation. Statistical Methodology, 17(??):67–81, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000640.

Hooten:2014:TVS

[469] M. B. Hooten, E. M. Hanks, D. S. Johnson, and M. W. Alldredge. Temporal variation and scale in movement-based resource selection functions. Statistical Methodology, 17(??):82–98, March 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000810.

Ghosh:2014:PMS

[470] Souparno Ghosh, David M. Bell, James S. Clark, Alan E. Gelfand, and Paul G. Flikkema. Process modeling for soil moisture using sensor network data. Statistical Methodology, 17(??):99–112, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000634.

Thurman:2014:VSS

[471] Andrew L. Thurman and Jun Zhu.

Variable selection for spatial Poisson point processes via a regularization method. Statistical Methodology, 17(??):113–125, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000622.

Leeds:2014:EAR

[472] W. B. Leeds, C. K. Wikle, and J. Fiechter. Emulator-assisted reducedrank ecological data assimilation for nonlinear multivariate dynamical spatiotemporal processes. Statistical Methodology, 17(??):126–138, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312712000792.

Chiu:2014:SSN

[473] Grace S. Chiu and Anton H. Westveld. A statistical social network model for consumption data in trophic food webs. Statistical Methodology, 17(??):139–160, March 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000713.

Anonymous:2014:EBb

[474] Anonymous. Editorial Board. Statistical Methodology, 17(??):IFC, March 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000877.

Tripathi:2014:IEP

[475] Yogesh Mani Tripathi, Somesh Kumar, and C. Petropoulos. Improved estimators for parameters of a Pareto distribution with a restricted scale. Statistical Methodology, 18(??):1–13, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000749.

Castro:2014:PLC

[476] Luis M. Castro, Victor H. Lachos, Guillermo P. Ferreira, and Reinaldo B. Arellano-Valle. Partially linear censored regression models using heavytailed distributions: a Bayesian approach. Statistical Methodology, 18 (??):14–31, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000853.

Franco-Pereira:2014:TTT

[477] Alba M. Franco-Pereira and Moshe Shaked. The total time on test transform and the decreasing percentile residual life aging notion. Statistical Methodology, 18(??):32–40, May 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000737.

Misra:2014:SBT

[478] Neeraj Misra and Mohd. Arshad. Selecting the best of two gamma populations having unequal shape parameters. Statistical Methodology, 18(??):41–63, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271300083X.

Chesneau:2014:NET

[479] Christophe Chesneau, Isha Dewan, and Hassan Doosti. Nonparametric estimation of a two dimensional continuous-discrete density function by wavelets. Statistical Methodology, 18(??):64–78, May 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000725.

Earls:2014:BCE

[480] Cecilia Earls and Giles Hooker.

Bayesian covariance estimation and inference in latent Gaussian process models. Statistical Methodology, 18 (??):79–100, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000828.

Dey:2014:GIE

[481] Sanku Dey and Biswabrata Pradhan. Generalized inverted exponential distribution under hybrid censoring. Statistical Methodology, 18 (??):101–114, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000701.

Raim:2014:MAF

[482] Andrew M. Raim, Minglei Liu, Nagaraj K. Neerchal, and Jorge G.

Morel. On the method of approximate Fisher scoring for finite mixtures of multinomials. Statistical Methodology, 18(??):115–130, May 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000841.

Ferreira:2014:BOS

[483] Marco A. R. Ferreira and Nilotpal Sanyal. Bayesian optimal sequential design for nonparametric regression via inhomogeneous evolutionary MCMC. Statistical Methodology, 18(??):131–141, May 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000981.

Anonymous:2014:EBc

[484] Anonymous. Editorial Board. Statistical Methodology, 18(??):IFC, May 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000033.

Mukhopadhyay:2014:IRW

[485] Nitis Mukhopadhyay. Independent ratios X/Z when Y/Z are (in)dependent but (X, Y), Z are dependent. Statistical Methodology, 19(??):1–3, July 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000112.

Bouzebda:2014:STP

[486] Salim Bouzebda and Khalid Chokri. Statistical tests in the partially linear additive regression models. Statistical Methodology, 19(??):4–24, July 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000161.

Akashi:2014:ELA

[487] Fumiya Akashi. Empirical likelihood approach toward discriminant analysis for dynamics of stable processes. Statistical Methodology, 19(??):25–43, July 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271400015X.

Jadhav:2014:SSM

[488] Nileshkumar H. Jadhav, Dattatraya N.

Kashid, and Subhash R. Kulkarni. Subset selection in multiple linear regression in the presence of outlier and multicollinearity. Statistical Methodology, 19(??):44–59, July 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000173.

Anonymous:2014:EBd

[489] Anonymous. Editorial Board. Statistical Methodology, 19(??):IFC, July 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000215.

Lahiri:2014:PPK

[490] Soumendra Lahiri and Min ge Xie.

Preface to the Professor Kesar Singh Memorial Issue. Statistical Methodology, 20(??):1, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000318.

Babu:2014:KSC

[491] G. Jogesh Babu. Kesar Singh’s contributions to statistical methodology. Statistical Methodology, 20 (??):2–10, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000993.

Bhattacharjee:2014:CLD

[492] Monika Bhattacharjee and Arup Bose. Consistency of large dimensional sample covariance matrix under weak dependence. Statistical Methodology, 20 (??):11–26, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000671.

Chakraborty:2014:DPD

[493] Anirvan Chakraborty and Probal Chaudhuri. The deepest point for distributions in infinite dimensional spaces. Statistical Methodology, 20 (??):27–39, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000427.

DasGupta:2014:SFB

[494] Anirban DasGupta, S. N. Lahiri, and Jordan Stoyanov. Sharp fixed n bounds and asymptotic expansions for the mean and the median of a Gaussian sample maximum, and applications to the Donoho–Jin model. Statistical Methodology, 20(??):40–62, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000124.

Datta:2014:BE

[495] Jyotishka Datta and Jayanta K.

Ghosh. Bootstrap — an exploration. Statistical Methodology, 20 (??):63–72, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000658.

Cohen:2014:MTI

[496] Arthur Cohen, Yingqiu Ma, Harold Sackrowitz, and Kesar Singh. Multiple testing and interval estimates of correlated correlations. Statistical Methodology, 20(??):73–78, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312713000440.

Serfling:2014:GFS

[497] Robert Serfling and Shanshan Wang.

General foundations for studying masking and swamping robustness of outlier identifiers. Statistical Methodology, 20(??):79–90, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271300066X.

Fourdrinier:2014:CGL

[498] Dominique Fourdrinier, William Strawderman, and Martin T. Wells. On completeness of the general linear model with spherically symmetric errors. Statistical Methodology, 20 (??):91–104, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312713000439.

Yang:2014:ENM

[499] Guang Yang, Dungang Liu, Regina Y. Liu, Minge Xie, and David C. Hoaglin. Efficient network metaanalysis: a confidence distribution approach. Statistical Methodology, 20 (??):105–125, September 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000136.

Anonymous:2014:EBe

[500] Anonymous. Editorial Board. Statistical Methodology, 20(??):IFC, September 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312714000331.

Beran:2014:EED

[501] Jan Beran and Georg Mainik. On estimating extremal dependence structures by parametric spectral measures. Statistical Methodology, 21 (??):1–22, November 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000185.

Barriga:2014:ZIC

[502] Gladys D. C. Barriga and Francisco Louzada. The zero-inflated Conway– Maxwell–Poisson distribution: Bayesian inference, regression modeling and influence diagnostic. Statistical Methodology, 21(??):23–34, November 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000148.

Ghosh:2014:BIN

[503] Anurag Ghosh, Soumalya Mukhopadhyay, Sandipan Roy, and Sourabh Bhattacharya. Bayesian inference in nonparametric dynamic state-space models. Statistical Methodology, 21 (??):35–48, November 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000197.

He:2014:EPD

[504] Hui He, Na Zhou, and Ruiming Zhang.

On estimation for the Pareto distribution. Statistical Methodology, 21 (??):49–58, November 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000288.

Bouzebda:2014:GTI

[505] Salim Bouzebda. General tests of independence based on empirical processes indexed by functions. Statistical Methodology, 21(??):59–87, November 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000276.

Economou:2014:KLD

[506] Polychronis Economou and George Tzavelas. Kullback–Leibler divergence measure based tests concerning the biasness in a sample. Statistical Methodology, 21(??):88–108, November 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000306.

Ng:2014:MSP

[507] Chi Tim Ng and Chi Wai Yu. Modified SCAD penalty for constrained variable selection problems. Statistical Methodology, 21(??):109–134, November 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000458.

Mukhopadhyay:2014:MCS

[508] Nitis Mukhopadhyay and Sankha Muthu Poruthotage. Multiple crossing sequential fixed-size confidence region methodologies for a multivariate normal mean vector. Statistical Methodology, 21 (??):135–155, November 2014. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271400029X.

Anonymous:2014:EBf

[509] Anonymous. Editorial Board. Statistical Methodology, 21(??):IFC, November 2014. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312714000501.

Kharrati-Kopaei:2015:NSC

[510] Mahmood Kharrati-Kopaei. A note on the simultaneous confidence intervals for the successive differences of exponential location parameters under heteroscedasticity. Statistical Methodology, 22(??):1–7, January 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271400046X.

Lee:2015:MET

[511] Jiyeon Lee, Sangyeol Lee, and Siyun Park. Maximum entropy test for GARCH models. Statistical Methodology, 22(??):8–16, January 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000471.

Mao:2015:CZE

[512] Chang Xuan Mao, Nan Yang, and Jinhua Zhong. On the Chao and Zelterman estimators in a binomial mixture model. Statistical Methodology, 22(??):17–22, January 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000483.

Nadarajah:2015:CDR

[513] Saralees Nadarajah, Sergey Bityukov, and Nikolai Krasnikov. Confidence distributions: a review. Statistical Methodology, 22(??):23–46, January 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000574.

Zhan:2015:IAI

[514] Xiang Zhan and Debashis Ghosh. Incorporating auxiliary information for improved prediction using combination of kernel machines. Statistical Methodology, 22(??):47–57, January 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000586.

Anonymous:2015:EBa

[515] Anonymous. Editorial Board. Statistical Methodology, 22(??):IFC, January 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000628.

Khosravi:2015:BMB

[516] Mohsen Khosravi, Debasis Kundu, and Ahad Jamalizadeh. On bivariate and a mixture of bivariate Birnbaum– Saunders distributions. Statistical Methodology, 23(??):1–17, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000562.

Cho:2015:ELI

[517] Youngseuk Cho, Hokeun Sun, and Kyeongjun Lee. Exact likelihood inference for an exponential parameter under generalized progressive hybrid censoring scheme. Statistical Methodology, 23(??):18–34, March 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000707.

Joekes:2015:EDS

[518] Silvia Joekes, Marcelo Smrekar, and Emanuel Pimentel Barbosa. Extending a double sampling control chart for non-conforming proportion in high quality processes to the case of small samples. Statistical Methodology, 23(??):35–49, March 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000719.

Chen:2015:ALF

[519] Tao Chen and Kenneth A. Couch. An approximation of logarithmic functions in the regression setting. Statistical Methodology, 23(??):50–58, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000720.

Klein:2015:MLE

[520] Daniel Klein and Ivan Zezula. Maximum likelihood estimators for extended growth curve model with orthogonal between-individual design matrices. Statistical Methodology, 23 (??):59–72, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000732.

Zhao:2015:SEL

[521] Peixin Zhao and Yiping Yang. Semiparametric empirical likelihood tests in varying coefficient partially linear models with repeated measurements. Statistical Methodology, 23 (??):73–87, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000768.

De:2015:STC

[522] Shyamal K. De and Michael Baron. Sequential tests controlling generalized familywise error rates. Statistical Methodology, 23(??):88–102, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

                                        http://www.sciencedirect.com/science/
                                        article/pii/S1572312714000744.

Han:2015:ESS

[523] David Han. Estimation in step-stress life tests with complementary risks from the exponentiated exponential distribution under time constraint and its applications to UAV data. Statistical Methodology, 23(??):103–122, March 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000604.

Anonymous:2015:EBb

[524] Anonymous. Editorial Board. Statistical Methodology, 23(??):IFC, March 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000793.

Wong:2015:DCH

[525] Tony Siu Tung Wong. Diagnostic check for heavy tail in linear time series. Statistical Methodology, 24(??):1–11, May 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S157231271400077X.

Satty:2015:DMH

[526] A. Satty, H. Mwambi, and G. Molenberghs. Different methods for handling incomplete longitudinal binary outcome due to missing at random dropout. Statistical Methodology, 24(??):12–27, May 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000756.

Rezapour:2015:SPS

[527] M. Rezapour and N. Balakrishnan.

Some properties of stochastic volatility model that are induced by its volatility sequence. Statistical Methodology, 24(??):28–36, May 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000884.

Khuri:2015:AMC

[528] A. I. Khuri, S. Mukhopadhyay, and M. A. Khuri. Approximating moments of continuous functions of random variables using Bernstein polynomials. Statistical Methodology, 24 (??):37–51, May 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000902.

Hossain:2015:SPP

[529] Shakhawat Hossain, S. Ejaz Ahmed, and Kjell A. Doksum. Shrinkage, pretest, and penalty estimators in generalized linear models. Statistical Methodology, 24(??):52–68, May 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312714000896.

Hosking:2015:URE

[530] J. R. M. Hosking and N. Balakrishnan. A uniqueness result for Lestimators, with applications to Lmoments. Statistical Methodology, 24(??):69–80, May 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312714000598.

Lee:2015:CPC

[531] Sangyeol Lee and Byungsoo Kim. Copula parameter change test for nonlinear AR models with nonlinear GARCH errors. Statistical Methodology, 25(??):1–22, July 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000027.

Kang:2015:FRC

[532] Dian tong Kang. Further results on closure properties of LPQE order. Statistical Methodology, 25 (??):23–35, July 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000052.

Chesneau:2015:SDA

[533] Christophe Chesneau, Maher Kachour, and Dimitris Karlis. On some distributions arising from a generalized trivariate reduction scheme. Statistical Methodology, 25(??):36–50, July 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000040.

Huang:2015:ATP

[534] Zhensheng Huang, Quanxi Shao, Zhen Pang, and Bingqing Lin. Adaptive testing for the partially linear single-index model with error-prone linear covariates. Statistical Methodology, 25(??):51–58, July 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000039.

Philip:2015:COS

[535] Anne Philip and P. Yageen Thomas.

On concomitants of order statistics arising from the extended Farlie– Gumbel–Morgenstern bivariate logistic distribution and its application in estimation. Statistical Methodology, 25(??):59–73, July 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000143.

Sudheesh:2015:SAT

[536] K. K. Sudheesh, P. Anisha, and C. M. Deemat. A simple approach for testing constant failure rate against different ageing classes for discrete data. Statistical Methodology, 25(??):74–88, July 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000131.

Fujisawa:2015:FSD

[537] Hironori Fujisawa and Toshihiro Abe.

A family of skew distributions with mode-invariance through transformation of scale. Statistical Methodology, 25(??):89–98, July 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000155.

Chalmond:2015:MDS

[538] Bernard Chalmond. A macro-DAG structure based mixture model. Statistical Methodology, 25(??):99– 118, July 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S1572312715000167.

Anonymous:2015:EBc

[539] Anonymous. Editorial Board. Statistical Methodology, 25(??):ifc, July 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000222.

Wang:2015:NPR

[540] Mingqiu Wang. Nonconvex penalized ridge estimations for partially linear additive models in ultrahigh dimension. Statistical Methodology, 26 (??):1–15, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000180.

Nourmohammadi:2015:DFT

[541] Mohammad Nourmohammadi, Mohammad Jafari Jozani, and Brad C. Johnson. Distribution-free tolerance intervals with nomination samples: Applications to mercury contamination in fish. Statistical Methodology, 26 (??):16–33, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000209.

Mukhopadhyay:2015:SNB

[542] Nitis Mukhopadhyay and Swarnali Banerjee. Sequential negative binomial problems and statistical ecology: a selected review with new directions. Statistical Methodology, 26 (??):34–60, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000192.

Provost:2015:DLA

[543] Serge B. Provost and Hyung-Tae Ha. Differentiated logdensity approximants. Statistical Methodology, 26 (??):61–71, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000179.

Cooray:2015:SML

[544] Kahadawala Cooray. A study of moments and likelihood estimators of the odd Weibull distribution. Statistical Methodology, 26(??):72– 83, September 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000295.

Krishnakumar:2015:TCR

[545] Jaya Krishnakumar and David Neto.

Testing for the cointegration rank in threshold cointegrated systems with multiple cointegrating relationships. Statistical Methodology, 26 (??):84–102, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000301.

Guillou:2015:EPS

[546] Armelle Guillou, Stéphane Loisel, and Gilles Stupfler. Estimating the parameters of a seasonal Markov-modulated Poisson process. Statistical Methodology, 26(??):103–123, September 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000325.

Rao:2015:FFP

[547] B. L. S. Prakasa Rao. Filtered fractional Poisson processes. Statistical Methodology, 26(??):124–134, September 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000337.

Kozubowski:2015:DTP

[548] Tomasz J. Kozubowski, Anna K. Panorska, and Matthew L. Forister. A discrete truncated Pareto distribution. Statistical Methodology, 26 (??):135–150, September 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000313.

Anonymous:2015:EB

[549] Anonymous. Editorial Board. Statistical Methodology, 26(??):ifc, September 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S1572312715000398.

Ghoreishi:2015:DBA

[550] S. K. Ghoreishi and M. R. Meshkani.

Dynamic Bayesian analysis of generalized odds ratios assuming multivariate skew-normal distribution for the error terms in the system equation. Statistical Methodology, 27 (??):1–11, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000349.

Shirazi:2015:MWB

[551] Esmaeil Shirazi and Hassan Doosti.

Multivariate wavelet-based density estimation with size-biased data. Statistical Methodology, 27(??):12–19, November 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312715000350.

Gai:2015:SIP

[552] Yujie Gai, Jun Zhang, Gaorong Li, and Xinchao Luo. Statistical inference on partial linear additive models with distortion measurement errors. Statistical Methodology, 27 (??):20–38, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000374.

Nair:2015:MDS

[553] N. Unnikrishnan Nair and P. G.

Sankaran. Multivariate discrete scalar hazard rate. Statistical Methodology, 27(??):39–50, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000362.

Alih:2015:CBI

[554] Ekele Alih and Hong Choon Ong. Cluster-based L_{2} re-weighted regression. Statistical Methodology, 27 (??):51–81, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000477.

Mitani:2015:NCW

[555] Aya A. Mitani, Allison W. Kurian, Amar K. Das, and Manisha Desai. Navigating choices when applying multiple imputation in the presence of multi-level categorical interaction effects. Statistical Methodology, 27(??):82–99, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000489.

Bobotas:2015:SST

[556] Panayiotis Bobotas and Maria Kateri. The step-stress tampered failure rate model under interval monitoring. Statistical Methodology, 27 (??):100–122, November 2015. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000490.

Gui:2015:DAS

[557] Wenhao Gui and Meiping Xu. Double acceptance sampling plan based on truncated life tests for half exponential power distribution. Statistical Methodology, 27(??):123–131, November 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000519.

Rudas:2015:DCP

[558] Tamás Rudas. Directionally collapsible parameterizations of multivariate binary distributions. Statistical Methodology, 27(??):132–145, November 2015. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000507.

Nandram:2016:BPI

[559] Balgobin Nandram and Jiani Yin.

Bayesian predictive inference under a Dirichlet process with sensitivity to the normal baseline. Statistical Methodology, 28(??):1–17, January 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000520.

Mahata:2016:CEN

[560] Kaushik Mahata, Amit Mitra, and Sharmishtha Mitra. Consistency of M-estimators of nonlinear signal processing models. Statistical Methodology, 28(??):18–36, January 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000532.

Anonymous:2016:EBa

[561] Anonymous. Editorial Board. Statistical Methodology, 28(??):ifc, January 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000672.

Lin:2016:CSM

[562] Gwo Dong Lin, Chin-Diew Lai, and K. Govindaraju. Correlation structure of the Marshall–Olkin bivariate exponential distribution. Statistical Methodology, 29(??):1–9, March 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000647.

Kossler:2016:ATT

[563] W. Kössler and Narinder Kumar. An adaptive test for the two-sample scale problem where the common quantile may be different from the median. Statistical Methodology, 29 (??):10–17, March 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000635.

Dewanji:2016:RAC

[564] Anup Dewanji, P. G. Sankaran, Debasis Sengupta, and Bappa Karmakar. Regression analysis of competing risks data with general missing pattern in failure types. Statistical Methodology, 29(??):18–31, March 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312715000659.

Rakitzis:2016:TPG

[565] Athanasios C. Rakitzis, Philippe Castagliola, and Petros E. Maravelakis. A two-parameter general inflated Poisson distribution: Properties and applications. Statistical Methodology, 29(??):32–50, March 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271500074X.

Kayid:2016:CMR

[566] M. Kayid, S. Izadkhah, and H. Alhalees.

Combination of mean residual life order with reliability applications. Statistical Methodology, 29(??):51–69, March 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000738.

Gorny:2016:ELI

[567] Julian Górny and Erhard Cramer. Exact likelihood inference for exponential distributions under generalized progressive hybrid censoring schemes. Statistical Methodology, 29(??):70–94, March 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000751.

Anonymous:2016:EBb

[568] Anonymous. Editorial Board. Statistical Methodology, 29(??):ifc, March 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000787.

Anonymous:2016:EBc

[569] Anonymous. Editorial Board. Statistical Methodology, 30(??):ifc, May 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000083.

Almasi:2016:MZI

[570] Afshin Almasi, Mohammad Reza Eshraghian, Abbas Moghimbeigi, Abbas Rahimi, Kazem Mohammad, and Sadegh Fallahigilan. Multilevel zeroinflated generalized Poisson regression modeling for dispersed correlated count data. Statistical Methodology, 30(??):1–14, May 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000763.

Shen:2016:ATM

[571] Pao sheng Shen. Analysis of transformation models with doubly truncated data. Statistical Methodology, 30(??):15–30, May 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000866.

Vossoughi:2016:DFT

[572] Mehrdad Vossoughi, S. M. T. Ayatollahi, Mina Towhidi, and Seyyed Taghi Heydari. A distribution-free test of parallelism for two-sample repeated measurements. Statistical Methodology, 30(??):31–44, May 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312715000878.

Espendiller:2016:FAM

[573] Michael Espendiller and Maria Kateri. A family of association measures for 2 × 2 contingency tables based on the ϕ-divergence. Statistical Methodology, 30(??):45–61, May 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000022.

Barakat:2016:NDA

[574] H. M. Barakat and A. R. Omar. A note on domains of attraction of the limit laws of intermediate order statistics under power normalization. Statistical Methodology, 31(??):1–7, July 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000034.

Bourguignon:2016:NSI

[575] Marcelo Bourguignon and Klaus L. P.

Vasconcellos. A new skew integer valued time series process. Statistical Methodology, 31(??):8–19, July 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000046.

Avlogiaris:2016:TLH

[576] G. Avlogiaris, A. Micheas, and K. Zografos. On testing local hypotheses via local divergence. Statistical Methodology, 31(??):20–42, July 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000058.

Ghashim:2016:BLB

[577] Ehssan Ghashim, Éric Marchand, and William E. Strawderman. On a better lower bound for the frequentist probability of coverage of Bayesian credible intervals in restricted parameter spaces. Statistical Methodology, 31(??):43–57, July 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000174.

Tarvirdizade:2016:ESS

[578] Bahman Tarvirdizade and Mohammad Ahmadpour. Estimation of the stress–strength reliability for the two-parameter bathtub-shaped lifetime distribution based on upper record values. Statistical Methodology, 31(??):58–72, July 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000162.

Chen:2016:ACN

[579] Cathy W. S. Chen, Mike K. P. So, Jessica C. Li, and Songsak Sriboonchitta. Autoregressive conditional negative binomial model applied to over-dispersed time series of counts. Statistical Methodology, 31(??):73–90, July 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000204.

Anonymous:2016:EBd

[580] Anonymous. Editorial Board. Statistical Methodology, 31(??):ifc, July 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000241.

Riggs:2016:CEP

[581] Kent R. Riggs, Phil D. Young, and Dean M. Young. Confidence ellipsoids for the primary regression coefficients in two seemingly unrelated regression models. Statistical Methodology, 32(??):1–13, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271600006X.

Kang:2016:DCR

[582] Dian-Tong Kang and Lei Yan. On the dynamic cumulative residual quantile entropy ordering. Statistical Methodology, 32(??):14–35, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000198.

Singh:2016:BOC

[583] Satya Prakash Singh and Siuli Mukhopadhyay. Bayesian optimal cluster designs. Statistical Methodology, 32 (??):36–52, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716000216.

Lui:2016:NEP

[584] Kung-Jong Lui. Notes on estimation in Poisson frequency data under an incomplete block crossover design. Statistical Methodology, 32 (??):53–62, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716000186.

Sordo:2016:TVO

[585] Miguel A. Sordo, Marilia C. de Souza, and Alfonso Suárez-Llorens. Testing variability orderings by using Gini’s mean differences. Statistical Methodology, 32(??):63–76, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716000228.

Katzur:2016:HTW

[586] Alexander Katzur and Udo Kamps. Homogeneity testing via weighted affinity in multiparameter exponential families. Statistical Methodology, 32 (??):77–90, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300028.

Yin:2016:RBK

[587] Jingjing Yin, Yi Hao, Hani Samawi, and Haresh Rochani. Rank-based kernel estimation of the area under the ROC curve. Statistical Methodology, 32(??):91–106, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300016.

Larose:2016:LCA

[588] Chantal Larose, Ofer Harel, Katarzyna Kordas, and Dipak K. Dey. Latent class analysis of incomplete data via an entropy-based criterion. Statistical Methodology, 32 (??):107–121, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300041.

Qian:2016:SIV

[589] Yunyun Qian and Zhensheng Huang.

Statistical inference for a varyingcoefficient partially nonlinear model with measurement errors. Statistical Methodology, 32(??):122–130, September 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312716300089.

Nematollahi:2016:LSM

[590] N. Nematollahi, R. Farnoosh, and Z. Rahnamaei. Location-scale mixture of skew-elliptical distributions: Looking at the robust modeling. Statistical Methodology, 32(??):131–146, September 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300053.

Ng:2016:GBO

[591] Chi Tim Ng, Seungyoung Oh, and Youngjo Lee. Going beyond oracle property: Selection consistency and uniqueness of local solution of the generalized linear model. Statistical Methodology, 32(??):147–160, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300107.

Kang:2016:EEV

[592] Xiaojuan Kang and Tizheng Li. Efficient estimation of varying coefficient models with serially correlated errors. Statistical Methodology, 32 (??):161–184, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300090.

Dey:2016:EPP

[593] Sanku Dey, Sukhdev Singh, Yogesh Mani Tripathi, and A. Asgharzadeh. Estimation and prediction for a progressively censored generalized inverted exponential distribution. Statistical Methodology, 32 (??):185–202, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300119.

Djeundje:2016:SDS

[594] Viani A. Biatat Djeundje. Systematic deviation in smooth mixed models for multi-level longitudinal data. Statistical Methodology, 32 (??):203–217, September 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300077.

Kang:2016:SNR

[595] Dian tong Kang. Some new results on the LQE ordering. Statistical Methodology, 32(??):218–235, September 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300120.

Anonymous:2016:EBe

[596] Anonymous. Editorial Board. Statistical Methodology, 32(??):ifc, September 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL

http://www.sciencedirect.com/science/
article/pii/S157231271630017X.

Kim:2016:HPA

[597] Sun-Hee Kim, Lei Li, Christos Faloutsos, Hyung-Jeong Yang, and Seong-Whan Lee. HeartCast: Predicting acute hypotensive episodes in intensive care units. Statistical Methodology, 33(??):1–13, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300132.

Stiburek:2016:STH

[598] David Stiburek. Sequential testing of hypotheses about drift for Gaussian diffusions. Statistical Methodology, 33(??):14–30, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300144.

Ceyhan:2016:EDN

[599] Elvan Ceyhan. Edge density of new graph types based on a random digraph family. Statistical Methodology, 33(??):31–54, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300156.

Yan:2016:SNR

[600] Lei Yan and Dian tong Kang. Some new results on the Rényi quantile entropy ordering. Statistical Methodology, 33(??):55–70, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271630003X.

Holte:2016:SDF

[601] Sarah E. Holte, Eva K. Lee, and Yajun Mei. Symmetric directional false discovery rate control. Statistical Methodology, 33(??):71–82, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300247.

Bustamante:2016:EGF

[602] Juan Carlos Bustamante and Edixon Chacón. Estimation and goodnessof-fit in latent trait models: a comparison among theoretical approaches. Statistical Methodology, 33 (??):83–95, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300065.

Hilgert:2016:CDU

[603] Nadine Hilgert, Ghislain Verdier, and Jean-Pierre Vila. Change detection for uncertain autoregressive dynamic models through nonparametric estimation. Statistical Methodology, 33 (??):96–113, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300259.

Brimacombe:2016:LSC

[604] Michael Brimacombe. Large sample convergence diagnostics for likelihood based inference: Logistic regression. Statistical Methodology, 33 (??):114–130, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300235.

Ferreira:2016:NRM

[605] Clécio S. Ferreira and Víctor H. Lachos. Nonlinear regression models under skew scale mixtures of normal distributions. Statistical Methodology, 33 (??):131–146, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300260.

Das:2016:DTS

[606] Sudipta Das, Anup Dewanji, and Debasis Sengupta. Discrete time software reliability modeling with periodic debugging schedule. Statistical Methodology, 33(??):147–159, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300284.

Almendra-Arao:2016:CTC

[607] Félix Almendra-Arao, José Juan Castro-Alva, and Hortensia Reyes-Cervantes. Constructing tests to compare two proportions whose critical regions guarantee to be Barnard convex sets. Statistical Methodology, 33 (??):160–171, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300272.

Luo:2016:FSE

[608] Shikai Luo and Subhashis Ghosal. Forward selection and estimation in high dimensional single index models. Statistical Methodology, 33(??): 172–179, December 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300302.

Zhan:2016:NPB

[609] Xiang Zhan and Debashis Ghosh. A novel power-based approach to Gaussian kernel selection in the kernel-based association test. Statistical Methodology, 33(??):180–191, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300314.

Khan:2016:EIM

[610] Rasul A. Khan. Estimating the integer mean of a normal model related to binomial distribution. Statistical Methodology, 33(??):192–202, December 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300326.

Wilding:2016:IPA

[611] Gregory E. Wilding and Mark C. Baker. Inference procedures about population correlations under order restrictions. Statistical Methodology, 33 (??):203–216, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300296.

Sim:2016:GIT

[612] Shin Zhu Sim and Seng Huat Ong.

A generalized inverse trinomial distribution with application. Statistical Methodology, 33(??):217–233, December 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S1572312716300338.

Chaouch:2016:NER

[613] Mohamed Chaouch, Naâmane Laïb, and Elias Ould Saïd. Nonparametric M-estimation for right censored regression model with stationary ergodic data. Statistical Methodology, 33(??):234–255, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S157231271630034X.

Bathaee:2016:NPB

[614] Najmeh Bathaee and Hamid Sheikhzadeh.

Non-parametric Bayesian inference for continuous density hidden Markov mixture model. Statistical Methodology, 33(??):256–275, December 2016. CO-DEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1572312716300351.

Anonymous:2016:EBf

[615] Anonymous. Editorial Board. Statistical Methodology, 33(??):ifc, December 2016. CODEN ???? ISSN 1572-3127 (print), 1878-0954 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1572312716300387.